### v1

# Spatially covariant gravity with two degrees of freedom in the presence of an auxiliary scalar field: Hamiltonian analysis

Abstract
A class of gravity theories respecting spatial covariance and in the presence of non-dynamical auxiliary scalar fields with only spatial derivatives is investigated. Generally, without higher temporal derivatives in the metric sector, there are 3 degrees of freedom (DOFs) propagating due to the breaking of general covariance. Through a Hamiltonian constraint analysis, we examine the conditions to eliminate the scalar DOF such that only 2 DOFs, which correspond the tensorial gravitational waves in a homogeneous and isotropic background, are propagating. We find that two conditions are needed, each of which can eliminate half degree of freedom. The second condition can be further classified into two cases according to its effect on the Dirac matrix. We also apply the formal conditions to a polynomial-type Lagrangian as a concrete example, in which all the monomials are spatially covariant scalars containing two derivatives. Our results are consistent with the previous analysis based on the perturbative method.
I Introduction
With the detection of gravitational waves, a novel tool has emerged for testing theories of gravity. A fundamental question arises: Is general relativity the only theory of gravity that propagates gravitational waves with precisely two degrees of freedom (2DOFs)? This question is partially addressed by Lovelock’s theorem [1, 2], which states that in four dimensions, if one requires spacetime covariance and second-order equations of motion, general relativity is indeed the unique theory that propagates gravitational waves with two degrees of freedom. This also implies that, by violating the assumptions of Lovelock’s theorem, other gravitational theories propagating two degrees of freedom can, in principle, exist.
Gravitational theories with 2DOFs can be traced back to the Cuscuton theory [3], which has been extensively studied [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], and extended in the framework of scalar-tensor theories with higher-order derivatives [17, 18]. Such theories have also been discussed as special cases of Hořava gravity [19, 20, 21, 22]. The connection between 2DOFs and spacetime symmetries has been investigated in [23, 24, 25, 26].
Another class of theories propagating 2DOFs was proposed in [27], known as minimally modified gravity (MMG) [^1] See also the minimal theory of massive gravity studied in [28, 29, 30, 31, 32, 33].. The so-called type-II MMG [34, 35] includes the original Cuscuton theory as a special case. MMG has been extensively studied [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], and has also been generalized by introducing auxiliary constraints in phase space [52, 53].
A general framework for spatially covariant gravity (SCG), which respects only spatial covariance, was proposed in [54, 55]. Due to the breaking of full spacetime covariance, SCG generally propagates a scalar degree of freedom in addition to the 2DOFs corresponding to gravitational waves. Thus, it provides an alternative approach to constructing scalar-tensor theories [56, 57, 58, 59]. SCG includes Hořava gravity [60, 61, 62] and the effective field theory of inflation/dark energy [63, 64, 65, 66, 67] as special cases. It has been further generalized to include a dynamical lapse function [68, 69], nonmetricity [70], parity violation [71], and multiple scalar modes [72]. SCG has also been applied in cosmology and gravitational wave studies [73, 74, 75, 76, 77]. Subclasses of SCG that propagate only 2DOFs were explored in [78, 79], and in the presence of a dynamical lapse function in [80]. The explicit Lagrangian found in [78] has been employed in cosmological studies [81, 82, 83, 10, 84].
In this work, we generalize the SCG framework to include an auxiliary scalar field and determine the conditions under which the theory propagates only 2DOFs. The idea of spatially covariant gravity with an auxiliary scalar field was first introduced in [85], originally motivated by generally covariant scalar-tensor theories in which the scalar field has a spacelike gradient [^2] However, we emphasize that these are entirely different theories. See Appendix A for details.. The presence of an auxiliary scalar field introduces new and novel possibilities for constructing 2DOF theories.
In [86], a perturbative analysis similar to [17, 69, 79] was performed to derive necessary conditions for the absence of scalar modes at linear order in perturbations around a cosmological background. Nevertheless, a Hamiltonian constraint analysis is still required to determine the necessary and sufficient conditions for the theory to propagate only 2DOFs at the nonlinear level. This paper is devoted to this task.
The paper is organized as follows. In Sec. II, we describe our model of spatially covariant gravity with an auxiliary scalar field. We find that new terms in the Lagrangian are allowed due to the inclusion of the auxiliary scalar. In Sec. III, we derive the Hamiltonian formalism and show that, in general, a scalar mode propagates unless additional constraints are imposed. In Sec. IV, we perform a degeneracy analysis and derive the conditions under which the theory propagates only 2DOFs. In Sec. V, we apply our formalism to a concrete model, where the Lagrangian is built from SCG monomials with $d=2$. Finally, in Sec. VI, we summarize our results.
II Spatially covariant gravity with an auxiliary scalar field
Our starting point is the action
$S=\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij};\mathrm{D}_{i},\pounds_{\bm{n}}\right),$ (1)
where $N$ is the usual lapse function, $h_{ij}$ is the spatial metric on the spacelike hypersurfaces, ${}^{3}\!R_{ij}$ is the spatial Ricci tensor, $\mathrm{D}_{i}$ is the spatially covariant derivative compatible with $h_{ij}$, and £$\pounds_{\bm{n}}$ is the Lie derivative with respect to the normal vector $n^{a}$ of the hypersurfaces. By definition, the action (1) respects only the three-dimensional spatial invariance. We require that the Lie derivative £$\pounds_{\bm{n}}$ acts only on $h_{ij}$, which yields the extrinsic curvature $K_{ij}$ defined by
£$\pounds_{\bm{n}}h_{ij}=2K_{ij}=\frac{1}{N}\left(\dot{h}_{ij}-\pounds_{\vec{N}}h_{ij}\right).$ (2)
As a result, both the lapse function $N$ and the scalar field $\phi$ play the role of auxiliary variables. Note that the shift vector $N^{i}$ should not appear explicitly in the Lagrangian, as it is actually the gauge field of spatial covariance.
Since the Lie derivative enters the Lagrangian only through the extrinsic curvature, we simply introduce $K_{ij}$ to replace the Lie derivative £$\pounds_{\bm{n}}$ in (1), and consider the following action:
$\tilde{S}=\int\mathrm{d}t\mathrm{d}^{3}x\,N\sqrt{h}\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right).$ (3)
According to the Hamiltonian analysis in the next section, the action (3) generally has 3 DOFs. This is simply because $\phi$ has no dynamics; the number of DOFs in (3) is the same as in spatially covariant gravity without the scalar field [54, 55].
In principle, Lie derivatives of the lapse function £$\pounds_{\bm{n}}N$ and of the extrinsic curvature £$\pounds_{\bm{n}}K_{ij}$ can be considered, which would introduce additional degrees of freedom. It has been shown that extra conditions must be imposed in order to eliminate these unwanted degrees of freedom. For example, the theory will propagate an extra DOF (i.e., a total of 4 DOFs) in the case of a dynamical lapse function. Two conditions are required to eliminate such an extra DOF [68]. A similar analysis was carried out to fully eliminate the scalar DOFs [80].
Generally, the Lagrangian in (3) may be nonlinear in $K_{ij}$, which makes the explicit inversion of velocity $\dot{h}_{ij}$ in terms of the momentum $\pi^{ij}$ impossible. This problem can be solved by rewriting (3) in an equivalent form:
$\tilde{S}$ $=\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\tilde{\mathcal{L}}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right)$
$=S\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\varLambda^{ij}\left(K_{ij}-B_{ij}\right),$ (4)
where the action $S\left(\phi,N,h_{ij},{}^{3}R_{ij},B_{ij};\mathrm{D}_{i}\right)$ is the same as (3) with $K_{ij}$ replaced by $B_{ij}$. This strategy of using a Lagrange multiplier to linearize the velocities was employed in [68, 78]. The equations of motion for $\varLambda^{ij}$ ensure that the auxiliary field $B_{ij}$ is exactly equal to $K_{ij}$. Thus, the two actions are equivalent, at least at the classical level. By making use of the equation of motion for the auxiliary field $B_{ij}$, we can determine the Lagrange multiplier $\varLambda^{ij}$ as
$\varLambda^{ij}=\frac{1}{N\sqrt{h}}\frac{\delta S}{\delta B_{ij}}.$ (5)
Finally, by plugging the solution for $\varLambda^{ij}$ into the action (4), we can recast it in the form
$\tilde{S}=S\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}x\frac{\delta S}{\delta B_{ij}}\left(K_{ij}-B_{ij}\right).$ (6)
In the following, we will use (6) as our starting point for the Hamiltonian analysis.
III Hamiltonian formalism
III.1 Hamiltonian and primary constraints
In the action (6), the set of variables is
$\varPhi^{I}\coloneqq\left\{N^{i},\phi,B_{ij},N,h_{ij}\right\},$ (7)
and the corresponding set of conjugate momenta is denoted by
$\varPi_{I}\coloneqq\left\{\pi_{i},p,p^{ij},\pi,\pi^{ij}\right\}.$ (8)
The conjugate momenta are defined as
$\varPi_{I}$ $\coloneqq\frac{\delta\tilde{S}}{\delta\dot{\Phi}^{I}}.$ (9)
Since all velocities $\dot{\varPhi}^{I}$ in the Lagrangian (6) appear linearly, the conjugate momenta automatically yield constraints among the phase space variables. Indeed, the conjugate momenta are
$\left\{\pi_{i},p,p^{ij},\pi,\pi^{ij}\right\}$ $=$ $\left\{\frac{\delta\tilde{S}}{\delta\dot{N}^{i}},\frac{\delta\tilde{S}}{\delta\dot{\phi}},\frac{\delta\tilde{S}}{\delta\dot{B}_{ij}},\frac{\delta\tilde{S}}{\delta\dot{N}},\frac{\delta\tilde{S}}{\delta\dot{h}_{ij}}\right\}$ (10)
$=$ $\left\{0^{i},0,0_{ij},0,\frac{1}{2N}\frac{\delta S}{\delta B_{ij}}\right\},$
which lead to the primary constraints
$\pi_{i}\approx 0,\quad p\approx 0,\quad p^{ij}\approx 0,\quad\pi\approx 0,\quad\tilde{\pi}^{ij}\equiv\pi^{ij}-\frac{1}{2N}\frac{\delta S}{\delta B_{ij}}\approx 0.$ (11)
Here and throughout this work, the symbol $\approx$” denotes “weak equality”, i.e., an equality that holds only when the constraints are satisfied. By contrast, “$=$” denotes strong equality, which holds regardless of whether constraints are imposed.
We denote the set of primary constraints as
$\varphi^{I}\coloneqq\left\{\pi_{i},p,p^{ij},\pi,\tilde{\pi}^{ij}\right\},$ (12)
and define the subspace of the phase space where the primary constraints are satisfied as $\varGamma_{\mathrm{P}}$, which can be regarded as a hypersurface in the original phase space $\left\{\varPhi^{I},\varPi_{I}\right\}$.
The canonical Hamiltonian is obtained via the Legendre transformation:
$H_{\mathrm{C}}$ $=\int\mathrm{d}^{3}x\left(\sum_{I}\varPi_{I}\dot{\varPhi}^{I}-N\sqrt{h}\tilde{\mathcal{L}}\right)\approx\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}},$ (13)
where
$\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}=\int\mathrm{d}^{3}x(NC)+X[\vec{N}],$ (14)
with
$C=2\pi^{ij}B_{ij}-\sqrt{h}\mathcal{L}.$ (15)
For a general “spatial” vector field $\vec{\xi}$, $X[\vec{\xi}]$ is defined as [68]
$X[\vec{\xi}]$$\coloneqq$ $\int\mathrm{d}^{3}x\,\varPi_{I}\pounds_{\vec{\xi}}\varPhi^{I}$ (16)
$=$ $\int\mathrm{d}^{3}x\left(\pi_{i}\pounds_{\vec{\xi}}N^{i}+p\pounds_{\vec{\xi}}\phi+p^{ij}\pounds_{\vec{\xi}}B_{ij}+\pi\pounds_{\vec{\xi}}N+\pi^{ij}\pounds_{\vec{\xi}}h_{ij}\right).$
In the following, we briefly summarize the properties of $X[\vec{\xi}]$ as discussed in [68]. Suppose $Q\left(\vec{x}\right)$ is a general spatial tensor field or density (with spatial indices suppressed for brevity) constructed from the phase space variables $\left\{\varPhi^{I},\varPi_{I}\right\}$, and let $\mathcal{F}$ be a scalar functional defined by
$\mathcal{F}\left[f,\varPhi^{I},\varPi_{I}\right]=\int\mathrm{d}^{3}x\,f(\vec{x})Q(\vec{x}),$ (17)
where $f(\vec{x})$ is a smooth test function. Note that $f(\vec{x})$ may also carry appropriate spatial indices (suppressed here) such that $\mathcal{F}$ is invariant under time-independent spatial diffeomorphisms. The Poisson bracket between $X[\vec{\xi}]$ and $\mathcal{F}$ is given by
$[X[\vec{\xi}],\mathcal{F}]$ $=$ $X\left[[\vec{\xi},\mathcal{F}]\right]+\int\mathrm{d}^{3}x\frac{\delta\mathcal{F}}{\delta f}\pounds_{\vec{\xi}}f$ (18)
$=$ $\int\mathrm{d}^{3}x\,Q(\vec{x})\pounds_{\vec{\xi}}f(\vec{x}).$
In (18), the Poisson bracket $[\mathcal{F},\mathcal{G}]$ is defined as
$[\mathcal{F},\mathcal{G}]\coloneqq\int\mathrm{d}^{3}z\left(\frac{\delta\mathcal{F}}{\delta\varPhi_{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPi^{I}(\vec{z})}-\frac{\delta\mathcal{F}}{\delta\varPi^{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPhi_{I}(\vec{z})}\right).$ (19)
A key point is that the right-hand side of the second line in (18) vanishes on the constraint surface if $Q(\vec{x})$ is a constraint.
By integrating by parts, $X[\vec{\xi}]$ can be written in a more convenient form:
$X[\vec{\xi}]=\int\mathrm{d}^{3}x\,\xi^{i}\mathcal{C}_{i},$ (20)
where
$\mathcal{C}_{i}$ $=$ $\pi\,\mathrm{D}_{i}N-2\sqrt{h}\mathrm{D}_{j}\left(\frac{1}{\sqrt{h}}\pi_{i}^{j}\right)+p\,\mathrm{D}_{i}\phi+p^{kl}\mathrm{D}_{i}B_{kl}$ (21)
$-2\sqrt{h}\mathrm{D}_{j}\left(\frac{p^{jk}}{\sqrt{h}}B_{ik}\right)+\pi_{j}\mathrm{D}_{i}N^{j}+\sqrt{h}\mathrm{D}_{j}\left(\frac{1}{\sqrt{h}}\pi_{i}N^{j}\right).$
As we shall see later, the constraints $\mathcal{C}_{i}\approx 0$ are secondary constraints associated with the primary constraints $\pi_{i}\approx 0$. Furthermore, the Poisson bracket between $\mathcal{C}_{i}$ and any constraint $Q$ is given by
$\left[X[\vec{\xi}],\mathcal{F}\right]=\int\mathrm{d}^{3}x\mathrm{d}^{3}y\,\xi^{i}(x)f(y)\left[\mathcal{C}_{i}(\vec{x}),Q(\vec{y})\right].$ (22)
It immediately follows from (18) that
$\left[\mathcal{C}_{i}(\vec{x}),Q(\vec{y})\right]\approx 0,\quad\text{for any}\quad Q\approx 0.$ (23)
Due to the presence of primary constraints, the time evolution is governed by the total Hamiltonian, defined as
$H_{\mathrm{T}}\coloneqq H_{\mathrm{C}}+\int\mathrm{d}^{3}y\sum_{I}\lambda_{I}(\vec{y})\varphi^{I}(\vec{y}),$ (24)
where $\lambda_{I}:=\left\{v^{i},v,v_{ij},\lambda,\lambda_{ij}\right\}$ are undetermined Lagrange multipliers. Since the canonical Hamiltonian $H_{\mathrm{C}}$ is only well-defined on the subspace defined by the primary constraints $\varGamma_{\mathrm{P}}$, we may directly use $\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}$ given in (14) in place of $H_{\mathrm{C}}$ in (13) for subsequent calculations.
III.2 Consistency condition
Constraints must be preserved under time evolution. For the primary constraints, we require that
$\int\mathrm{d}^{3}y\left[\varphi^{I}(\vec{x}),\varphi^{J}(\vec{y})\right]\lambda_{J}(\vec{y})+\left[\varphi^{I}(\vec{x}),H_{\mathrm{C}}\right]\approx 0,$ (25)
which are the so-called consistency conditions for the primary constraints. These conditions may yield further constraints.
Without imposing additional assumptions on the structure of the Lagrangian in (6), the Poisson brackets between the primary constraints have been computed in detail in [85], and are summarized in Appendix B. For later convenience, we rewrite (25) in matrix form:
$\int\mathrm{d}^{3}y\left(\begin{array}[]{ccccc}0&0&0&0&0\\ 0&0&0&0&\left[p(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&0&0&0&\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&0&0&0&\left[\pi(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&\left[\tilde{\pi}^{ij}(\vec{x}),p(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),p^{kl}(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),\pi(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\end{array}\right)\left(\begin{array}[]{c}v^{i}\\ v\\ v_{ij}\\ \lambda\\ \lambda_{ij}\end{array}\right)\approx\left(\begin{array}[]{c}-\mathcal{C}_{i}(\vec{x})\\ \frac{\delta S}{\delta\phi(\vec{x})}\\ 0\\ -\mathcal{C}(\vec{x})\\ \left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]\end{array}\right),$ (26)
where $\mathcal{C}_{i}$ is defined in (22). Note that $\mathcal{C}_{i}$ must have vanishing Poisson brackets with all other constraints due to the property (22). In the fourth line of (26), $\mathcal{C}$ is defined by
$\mathcal{C}=2\pi^{ij}(\vec{x})B_{ij}(\vec{x})-\frac{\delta S}{\delta N(\vec{x})},$ (27)
and $\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]$ in the fifth line is given by
$\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]=\frac{\delta S}{\delta h_{ij}(\vec{x})}-\frac{1}{N(\vec{x})}\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{z})}N(\vec{z})B_{kl}(\vec{z}).$ (28)
According to (129), $\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ in the third line of (26) is proportional to the kinetic matrix
$\frac{\delta^{2}S_{B}}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})},$ (29)
which we assume to be non-degenerate. This is because $B_{ij}$ is the auxiliary field corresponding to $K_{ij}$. The non-degeneracy of (29) implies that the kinetic term for $\dot{h}_{ij}$ is non-degenerate, which guarantees that the theory includes General Relativity as a limiting case, and in particular, allows the existence of gravitational waves.
As a result, the Lagrange multipliers $\lambda_{ij}$ can be fixed from the third line of (26) as
$\lambda_{ij}=0.$ (30)
Then the second and fourth lines of (26) yield two additional secondary constraints:
$\frac{\delta S}{\delta\phi(\vec{x})}\approx 0,\quad\mathcal{C}(\vec{x})\approx 0.$ (31)
The constraints $\frac{\delta S}{\delta\phi(\vec{x})}$ and $\mathcal{C}(\vec{x})$ do not generate further constraints via consistency conditions, as will become clear in the next section.
Having identified all the primary and secondary constraints, we are now in a position to classify them. According to the terminology of the Dirac-Bergmann algorithm, a constraint that has vanishing Poisson brackets with all other constraints (including itself) is called a first-class constraint; otherwise, it is second-class. The distinction is that each first-class constraint eliminates one degree of freedom, whereas a second-class constraint removes one phase space dimension, i.e., half a degree of freedom.
From the discussion above on $\mathcal{C}_{i}$ and the constraint algebra in Appendix B, the Poisson brackets among all the constraints can be summarized in the so-called Dirac matrix:
$\begin{array}[]{|c|cccccccc|}\hline\cr[\cdot,\cdot]&\pi_{i}(\vec{y})&\mathcal{C}_{i}(\vec{y})&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\tilde{\pi}^{kl}(\vec{y})&\mathcal{C}(\vec{y})&\frac{\delta S}{\delta\phi(\vec{y})}\\ \hline\cr\pi_{i}(\vec{x})&0&0&0&0&0&0&0&0\\ \mathcal{C}_{i}(\vec{x})&0&0&0&0&0&0&0&0\\ p(\vec{x})&0&0&0&0&0&X&X&X\\ \pi(\vec{x})&0&0&0&0&0&X&X&X\\ p^{ij}(\vec{x})&0&0&0&0&0&X&X&X\\ \tilde{\pi}^{ij}(\vec{x})&0&0&X&X&X&X&X&X\\ \mathcal{C}(\vec{x})&0&0&X&X&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&0&0&X&X&X&X&X&0\\ \hline\cr\end{array},$ (32)
where an “$X$” denotes a generically non-vanishing Poisson bracket. In total, there are 22 constraints, which can be divided into two classes:
$\begin{array}[]{rl}6\text{ first-class: }&\pi_{i},\quad\mathcal{C}_{i},\\ 16\text{ second-class: }&p,\quad p^{ij},\quad\pi,\quad\tilde{\pi}^{ij},\quad\frac{\delta S}{\delta\phi},\quad\mathcal{C}.\end{array}$
The number of degrees of freedom (DOF) is thus
#$\#_{\mathrm{DOF}}$ $=$ $\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (33)
$=$ $\frac{1}{2}(2\times 17-2\times 6-16)$
$=$ 3 3,3 ,
which explicitly shows that a spatially covariant gravity theory where temporal derivatives arise only through the extrinsic curvature has 3 propagating degrees of freedom, regardless of the presence of additional auxiliary field(s).
IV Degeneracy Analysis
To reduce the number of DOFs, additional conditions must be imposed on the Lagrangian (6). As we will see in this section, two conditions are required in order for the theory to propagate only two DOFs.
IV.1 The first condition
As first-class constraints, $\pi_{i}$ and $\mathcal{C}_{i}$ correspond to spatial invariance. Without breaking spatial invariance, $\pi_{i}$ and $\mathcal{C}_{i}$ will always remain first-class constraints, which have vanishing Poisson brackets with all other constraints. For simplicity, we are allowed to omit these two first-class constraints in the subsequent degeneracy analysis. We therefore define a simplified Dirac matrix by omitting the two columns and rows corresponding to $\pi_{i}$ and $\mathcal{C}_{i}$:
$\mathcal{M}^{ab}(\vec{x},\vec{y})=\begin{array}[]{|c|ccc|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\tilde{\pi}^{kl}(\vec{y})&\mathcal{C}(\vec{y})&\frac{\delta S}{\delta\phi(\vec{y})}\\ \hline\cr p(\vec{x})&0&0&0&X&X&X\\ \pi(\vec{x})&0&0&0&X&X&X\\ p^{ij}(\vec{x})&0&0&0&X&X&X\\ \hline\cr\tilde{\pi}^{ij}(\vec{x})&X&X&X&X&X&X\\ \mathcal{C}(\vec{x})&X&X&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X&X&X&0\\ \hline\cr\end{array}.$ (34)
Our goal is to find conditions that eliminate one scalar-type degree of freedom. Generally, we have two options: one is to introduce extra conditions to convert second-class constraints into first-class constraints, and the other is to generate new secondary constraints. In both cases, a necessary condition is that the simplified Dirac matrix $\mathcal{M}^{ab}$ must be degenerate. The determinant of the simplified Dirac matrix $\mathcal{M}^{ab}$ is
$\det\mathcal{M}^{ab}(\vec{x},\vec{y})=\det\mathcal{B}^{ab}(\vec{x},\vec{y})\times\det\mathcal{B}^{ba}(\vec{y},\vec{x}),$ (35)
where the matrix $\mathcal{B}^{ab}$ is the lower-left corner of $\mathcal{M}^{ab}$. The degeneracy of $\mathcal{M}^{ab}$, i.e., $\det\mathcal{M}^{ab}=0$, thus implies that the sub-matrix $\mathcal{B}^{ab}$ of $\mathcal{M}^{ab}$ is degenerate, i.e., $\det\mathcal{B}^{ab}=0$.
At first glance, the determinant of the matrix $\mathcal{B}^{ab}$ may appear to be the matrix form of the first 2DOF condition. However, this is not the case. The 2DOF condition should be an equation satisfied by the action and should not contain conjugate momenta. Otherwise, we obtain a relation between the action and the conjugate momenta, which is in fact a constraint. Upon careful inspection, all elements of the matrix $\mathcal{B}^{ab}$ are independent of the conjugate momenta, except for $\mathcal{B}^{21}$, given by
$\mathcal{B}^{21}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}(\vec{x}),p^{kl}(\vec{y})\right]$ (36)
$=$ $2\delta^{3}(\vec{x}-\vec{y})\pi^{kl}(\vec{x})-\frac{\delta^{2}S}{\delta N(\vec{x})\delta B_{kl}(\vec{y})},$
which can be rewritten as
$\mathcal{B}^{21}(\vec{x},\vec{y})\approx-\delta^{3}(\vec{x}-\vec{y})\frac{1}{N(\vec{y})}\frac{\delta S}{\delta B_{ij}(\vec{y})}+\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta N(\vec{y})}$ (37)
on the constraint surface of $\tilde{\pi}^{ij}$. In subsequent calculations, we will always use (37) as the value of $\mathcal{B}^{21}$ instead of (36). As a result, the matrix form of the first 2DOF condition becomes
$\mathcal{S}_{1}(\vec{x},\vec{y})=\det\mathcal{B}^{ab}(\vec{x},\vec{y})\approx 0,$ (38)
where
$\mathcal{B}^{ab}=\begin{array}[]{|c|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})\\ \hline\cr-2N(\vec{x})\tilde{\pi}^{ij}(\vec{x})&\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})}&N(\vec{x})\frac{\delta}{\delta N(\vec{y})}\left(\frac{1}{N(\vec{x})}\frac{\delta S}{\delta B_{ij}(\vec{x})}\right)&\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})}\\ -\mathcal{C}(\vec{x})&\frac{\delta^{2}S}{\delta N(\vec{x})\delta\phi(\vec{y})}&\frac{\delta^{2}S}{\delta N(\vec{x})\delta N(\vec{y})}&N(\vec{y})\frac{\delta}{\delta N(\vec{x})}\left(\frac{1}{N(\vec{y})}\frac{\delta S}{\delta B_{kl}(\vec{y})}\right)\\ \frac{\delta S}{\delta\phi(\vec{x})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta N(\vec{y})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta B_{kl}(\vec{y})}\\ \hline\cr\end{array}.$ (39)
In the above, we have adjusted the coefficients of $\tilde{\pi}^{ij}$ and $\mathcal{C}$ in the matrix $\mathcal{B}^{ab}$ to make it more symmetric. Since $\mathcal{B}^{ab}$ is a degenerate matrix, there must exist a null eigenvector $\mathcal{V}_{a}$ satisfying
$\int\mathrm{d}^{3}y\,\mathcal{B}^{ab}(\vec{x},\vec{y})\mathcal{V}_{b}(\vec{y})\approx 0^{a}.$ (40)
Generally, $\mathcal{V}_{a}$ takes the form
$\mathcal{V}_{a}(\vec{x})=\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x})=\int\mathrm{d}^{3}y\,U_{a}(\vec{x},\vec{y})V(\vec{y}).$ (41)
In (41), $V(\vec{x})$ is an arbitrary function of the spatial coordinates, and $U_{a}(\vec{x},\vec{y})$ depends on the phase space variables. One method to obtain the explicit form of the null eigenvector $\mathcal{V}_{a}$ is to calculate the adjoint matrix of $\mathcal{B}^{ab}$, as detailed in Appendix C.
Assuming that $\mathcal{B}_{2}^{*}$ does not vanish, we can use a linear combination of the null eigenvector $\mathcal{V}_{a}$ and the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ to define a new constraint $\mathcal{C}^{\prime}$,
$\int\mathrm{d}^{3}x\,\mathcal{C}^{\prime}(\vec{x})V(\vec{x})=\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}\tilde{\pi}^{ij}&\mathcal{C}&\frac{\delta S}{\delta\phi}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x})\approx 0.$ (42)
In (42), $V(\vec{x})$ is the arbitrary function from (41).
The advantage of introducing the new constraint $\mathcal{C}^{\prime}$ is that its Poisson brackets with the constraints $p$, $\pi$, and $p^{kl}$ all vanish:
$\left[\mathcal{C}^{\prime}(\vec{x}),p(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),\pi(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),p^{kl}(\vec{y})\right]\approx 0.$ (43)
With this new set of primary constraints, the consistency condition for the new constraint $\mathcal{C}^{\prime}$ becomes
$\frac{\mathrm{d}\mathcal{C}^{\prime}(\vec{x})}{\mathrm{d}t}$ $=\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{T}}\right]$ (44)
$=\int\mathrm{d}^{3}y\left[\mathcal{C}^{\prime}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\lambda_{kl}(\vec{y})+\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\approx 0$
where $\lambda_{ij}$ has been fixed to zero in (26). If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$, (44) yields a new secondary constraint $\Phi(\vec{y})$:
$\Phi(\vec{y})=\left[\mathcal{C}^{\prime}(\vec{y}),H_{C}\right]\approx 0.$ (45)
Otherwise, if no secondary constraint arises, we can still find a new first-class constraint by taking a linear combination of the existing constraints $p$, $\pi$, and $p^{ij}$. The necessary and sufficient condition for generating secondary constraints is that the consistency condition is not automatically satisfied (instead of requiring the simplified Dirac matrix to be degenerate). Therefore, after imposing the consistency condition, two scenarios may arise:
  1. 1.
    The consistency condition is not automatically satisfied, and an additional secondary constraint is needed to enforce it;
  2. 2.
    The consistency condition is automatically satisfied, and a first-class constraint can be constructed from a linear combination of existing constraints.
When we impose a specific 2DOF condition, we must check whether it violates the consistency condition and leads to additional secondary constraints. Such a newly generated constraint must be taken into account in the subsequent analysis.
IV.2 The second condition
If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$ in the previous section, we introduce a new constraint $\pi^{\prime}$ in the same way as $\mathcal{C}^{\prime}$,
$\int\mathrm{d}^{3}x\,\mathcal{\pi}^{\prime}(\vec{x})V(\vec{x})\equiv\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}p^{ij}&\pi&p\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}^{\prime}_{1}\right)_{ij}&\mathcal{\mathcal{V}}^{\prime}_{2}&\mathcal{\mathcal{V}}^{\prime}_{3}\end{array}\right)(\vec{x})\approx 0,$ (46)
where again $V(\vec{x})$ is the arbitrary function in (41). Similar to the constraint $\mathcal{C}^{\prime}$, the new constraint $\mathcal{\pi}^{\prime}$ has vanishing Poisson brackets:
$\left[\tilde{\pi}^{ij}(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]\approx\left[\frac{\delta S}{\delta\phi(\vec{x})},\mathcal{\pi}^{\prime}(\vec{y})\right]\approx 0.$ (47)
By making use of the constraints $\mathcal{\pi}^{\prime}$, $\mathcal{C}^{\prime}$ and $\Phi$, the simplified Dirac matrix becomes
=$\mathcal{M}^{\prime}{}^{ab}=\begin{array}[]{|c|ccccc|}\hline\cr[\cdot,\cdot]&\mathcal{\pi}^{\prime}(\vec{y})&\left\{p(\vec{y}),p^{kl}(\vec{y})\right\}&\mathcal{C}^{\prime}(\vec{y})&\left\{\tilde{\pi}^{kl}(\vec{y}),\frac{\delta S}{\delta\phi(\vec{y})}\right\}&\Phi(\vec{x})\\ \hline\cr\mathcal{\pi}^{\prime}(\vec{x})&0&0&0&0&X\\ \left\{p(\vec{x}),p^{kl}(\vec{x})\right\}&0&0&0&-\mathcal{C}^{ba}(\vec{y},\vec{x})&X\\ \mathcal{C}^{\prime}(\vec{x})&0&0&X&X&X\\ \left\{\tilde{\pi}^{kl}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{x})}\right\}&0&\mathcal{C}^{ab}(\vec{x},\vec{y})&X&X&X\\ \Phi(\vec{x})&X&X&X&X&X\\ \hline\cr\end{array}$ (48)
Since the first 2DOF condition $\mathcal{S}_{1}(\vec{x},\vec{y})$ has been imposed, which yields a secondary constraint that reduces the dimension of the phase space by one and thus removes half a degree of freedom, we need to impose an additional 2DOF condition to reduce the number of degrees of freedom to two. This means that the simplified Dirac matrix must be degenerate again. The determinant of this matrix is proportional to $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]$, $\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]$, and $\det\mathcal{C}^{ab}(\vec{x},\vec{y})$. If the determinant of this matrix is required to vanish, two types of 2DOF conditions can be obtained:
$\mathcal{S}_{2}^{(1)}(\vec{x},\vec{y})$ $=$ $\det\mathcal{C}^{ab}(\vec{x},\vec{y})\times\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]=\det\mathcal{D}^{ab}(\vec{x},\vec{y})\approx 0,$ (49)
$\mathcal{S}_{2}^{(2)}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]\approx 0,$ (50)
where
$\mathcal{D}^{ab}(\vec{x},\vec{y})$ $=\begin{array}[]{|c|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})\\ \hline\cr\tilde{\pi}^{ij}(\vec{x})&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X\\ \Phi(\vec{x})&X&X&X\\ \hline\cr\end{array}.$ (55)
The 2DOF condition $\mathcal{S}_{2}^{(1)}$ will generate a secondary constraint if the consistency condition is not automatically satisfied. On the other hand, the 2DOF condition $\mathcal{S}_{2}^{(2)}$ clearly generates a first-class constraint.
If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]=0$ in the previous section, the second 2DOF conditions become
$\mathcal{S}^{\prime}{}_{2}^{(1)}(\vec{x},\vec{y})$ $=$ $\det\mathcal{C}^{ab}(\vec{x},\vec{y})\approx 0,$ (56)
$\mathcal{S}^{\prime}{}_{2}^{(2)}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]\approx 0,$ (57)
which are covered by the TTDOF conditions $\mathcal{S}_{2}^{(1)}$ and $\mathcal{S}_{2}^{(2)}$.
The space of theories satisfying the 2DOF conditions $\mathcal{S}_{1}$, $\mathcal{S}{}_{2}^{(1)}$, and $\mathcal{S}_{2}^{(2)}$ can be divided into three branches:
  1. 1.
    Linear dependence branch: If the secondary constraints are linearly dependent, the 2DOF conditions will be automatically satisfied. However, in this case, the number of degrees of freedom will not change. Due to the linear dependence, the number of secondary constraints is reduced by one, which offsets the half-degree-of-freedom reduction caused by the 2DOF conditions. For example, if the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ are linearly dependent, the 2DOF condition $\mathcal{S}_{1}$ is automatically satisfied. However, the number of degrees of freedom remains unchanged.
  2. 2.
    Non-physical branch: Some 2DOF conditions may render the theory trivial. For example, if the secondary constraints generated by the consistency condition still fail to satisfy the consistency condition, additional secondary constraints will be generated. These additional constraints may reduce the number of degrees of freedom to less than two, which is unphysical.
  3. 3.
    Physical branch: In this branch, we obtain a scalar-tensor theory with only two tensorial degrees of freedom.
In the following, when we refer to 2DOF conditions, we always mean those corresponding to the third physical branch.
V A concrete model of $d=2$
V.1 The Lagrangian
In [57], [58] (see also [74]), monomials of SCG have been constructed and classified according to the total number of derivatives $d$ in each monomial. As an example of our formalism, in this section we apply the two conditions to the concrete model with $d=2$. The action is given by
$\tilde{S}_{2}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right)=S_{2}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}x\frac{\delta S_{2}}{\delta B_{ij}}\left(K_{ij}-B_{ij}\right).$ (58)
As discussed earlier, we introduce the variable $B_{ij}$ to make the explicit inversion of the velocity $\dot{h}_{ij}$ in terms of the momentum possible. The equivalent action $S_{2}$ can be written as
$S_{2}=\int\mathrm{d}t\mathrm{d}^{3}x\,N\sqrt{h}\mathcal{L}_{2},$ (59)
where
$\mathcal{L}_{2}=c_{1}B_{ij}B^{ij}+c_{2}B^{2}+c_{3}R+c_{4}a_{i}a^{i}+d_{1}\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi+d_{2}a_{i}\mathrm{D}^{i}\phi.$ (60)
The coefficients $c_{i}$ and $d_{i}$ are general functions of $N$ and $\phi$. The acceleration $a_{i}$ is defined as
$a_{i}=\frac{1}{N}\mathrm{D}_{i}N.$ (61)
For later convenience, we evaluate some of the elements of $\mathcal{B}^{ab}$ as
$\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{mn}(\vec{y})}$ $=$ $2N\sqrt{h}\left(\frac{1}{2}c_{1}(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm})+c_{2}\delta^{ij}\delta^{mn}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}),$ (62)
$N(\vec{y})\frac{\delta}{\delta N(\vec{x})}\left(\frac{1}{N(\vec{y})}\frac{\delta S_{B}}{\delta B_{ij}(\vec{y})}\right)$ $=$ $2N\sqrt{h}\left(\frac{\partial c_{1}}{\partial N}B^{ij}+\frac{\partial c_{2}}{\partial N}B\delta^{ij}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}),$ (63)
$\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})}$ $=$ $2N\sqrt{h}\left(\frac{\delta c_{1}}{\delta\phi}B^{ij}+\frac{\delta c_{2}}{\delta\phi}B\delta^{ij}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}).$ (64)
To prevent the dynamical terms in the theory from degenerating, we assume
$\det\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{mn}(\vec{y})}=\left(2N\sqrt{h}\delta^{3}(\vec{x}-\vec{y})\right)^{6}c_{1}^{6}(c_{1}+3c_{2})\neq 0,$ (65)
which implies
$c_{1}\neq 0,\qquad c_{1}+3c_{2}\neq 0.$ (66)
As long as (66) is satisfied, the non-physical branch of the 2DOF conditions is naturally removed.
Other elements of the matrix $\mathcal{B}^{ab}$ are evaluated as
$\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}+\right.$ (67)
$\quad\left(-2\frac{\partial d_{1}}{\partial\phi}\mathrm{D}^{i}\phi+\Big{(}-2N\frac{\partial d_{1}}{\partial N}-2d_{1}\Big{)}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-2d_{1}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
$\frac{\delta^{2}S}{\delta N(\vec{x})\delta N(\vec{y})}$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta N\delta N}\right)_{0}+\right.$ (68)
$\quad+\left(-\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathrm{D}^{i}\phi+\Big{(}\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\Big{)}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-\frac{2c_{4}}{N^{2}}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
$\frac{\delta}{\delta\phi(\vec{y})}\left(\frac{\delta S}{\delta N(\vec{x})}\right)$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta\phi\delta N}\right)_{0}+\right.$ (69)
$\quad\left(\Big{(}\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\Big{)}\mathrm{D}^{i}\phi-\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-\frac{d_{2}}{N}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
$\frac{\delta}{\delta N(\vec{y})}\left(\frac{\delta S}{\delta\phi(\vec{x})}\right)$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta N\delta\phi}\right)_{0}+\right.$ (70)
$\quad\left(\Big{(}-2\frac{\partial d_{1}}{\partial N}-\frac{2d_{1}}{N}\Big{)}\mathrm{D}^{i}\phi+\Big{(}\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\Big{)}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-\frac{d_{2}}{N}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
where $\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}$ represents the terms in $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ that are independent of derivatives of the delta function. We use this notation because their explicit expressions are very complicated, and only the coefficients of the derivatives of the delta function are relevant to the subsequent analysis.
V.2 The first condition
We have already obtained the first 2DOF condition in (38). To get the specific form of the 2DOF conditions for this model, a straightforward approach is to calculate the determinant of the matrix $\mathcal{B}^{ab}$, which is, however, very complicated and not illuminating. Alternatively, we can choose an equivalent approach by assuming that a null eigenvector $\mathcal{V}_{a}(\vec{x})$ exists and find the expression of this null eigenvector. In this approach, we can determine which conditions are required to make the Dirac matrix $\mathcal{B}^{ab}$ degenerate.
We assume $U_{a}(\vec{x},\vec{y})$ in (41) is independent of the derivative of the delta function. As a result, different orders of derivatives of the delta function will be separated in the matrix $\mathcal{B}^{ab}$,
$\mathcal{B}^{ab}(\vec{x},\vec{y})=\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})+\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})+\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y}),$ (71)
in which terms with different orders (of derivatives of the delta function) in the matrix $\mathcal{B}^{ab}$ separately satisfy
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(i)}^{ab}(\vec{x},\vec{y})\approx 0^{b},\quad i=0,1,2.$ (72)
Note that by observing the form of the secondary constraint $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$, (72) can only be weakly satisfied when $i=0$.
In the matrices $\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})$ and $\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y})$, only the $\times 2$ 2 submatrices in the lower left corner are nonvanishing, which are
$\left(\begin{array}[]{cc}2\left(\frac{d_{1}}{N}+\frac{\partial d_{1}}{\partial N}-\frac{1}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathrm{D}^{i}\phi-\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}a^{i}&-\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathrm{D}^{i}\phi+2\left(\frac{c_{4}}{N^{2}}-\frac{1}{N}\frac{\partial c_{4}}{\partial N}\right)a^{i}\\ -2\frac{\partial d_{1}}{\partial\phi}\mathrm{D}^{i}\phi-2\left(N\frac{\partial d_{1}}{\partial N}+d_{1}\right)a^{i}&-2\left(\frac{\partial d_{1}}{\partial N}+\frac{d_{1}}{N}\right)\mathrm{D}^{i}\phi+\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)a^{i}\end{array}\right)_{x}\mathrm{D}_{x^{i}}\delta^{3}\left(\vec{x}-\vec{y}\right)$ (73)
and
$\left(\begin{array}[]{cc}-\frac{d_{2}}{N}&-\frac{2c_{4}}{N^{2}}\\ -2d_{1}&-\frac{d_{2}}{N}\end{array}\right)_{x}\mathrm{D}^{x^{i}}\mathrm{D}_{x_{i}}\delta^{3}\left(\vec{x}-\vec{y}\right),$ (74)
respectively. In the above, a subscript “$x$” denotes functional dependence on $\vec{x}$. In the subsequent discussion, we omit $\vec{x}$ for brevity when the dependence on $\vec{x}$ is clear from the context.
According to (74), for the null eigenvalue equation of $\mathcal{B}_{(2)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ proportional to $\mathrm{D}^{x^{i}}\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
$d_{2}\mathcal{V}_{2}+2Nd_{1}\mathcal{V}_{3}$ $=$ 0 0,0 , (75)
$2c_{4}\mathcal{V}_{2}+Nd_{2}\mathcal{V}_{3}$ $=$ 0 0.0 . (76)
As for the null eigenvalue equation of $\mathcal{B}_{(1)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ only related to $\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
$\left(\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathcal{V}_{2}-2\frac{\partial d_{1}}{\partial\phi}\mathcal{V}_{3}$ $=$ 0 0,0 , (77)
$\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2N\frac{\partial d_{1}}{\partial N}+2d_{1}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (78)
$\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2\frac{\partial d_{1}}{\partial N}+\frac{2d_{1}}{N}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (79)
$\left(\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\right)\mathcal{V}_{2}-\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)\mathcal{V}_{3}$ $=$ 0 0.0 . (80)
There are four equations since the coefficients of $\mathrm{D}^{i}\phi$ and $a^{i}$ need to be proportional to the null eigenvector. We assume that $d_{1}$, $d_{2}$, and $c_{4}$ are nonvanishing. We can simply choose
$\mathcal{V}_{2}$ $=$ $2Nd_{1},$ (81)
$\mathcal{V}_{3}$ $=$ $-d_{2},$ (82)
and obtain two independent equations
$d_{2}^{2}-4d_{1}c_{4}$ $=$ 0 0,0 , (83)
$\left(\frac{d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)2Nd_{1}+\frac{\partial d_{1}}{\partial\phi}d_{2}$ $=$ 0 0.0 . (84)
The other two equations are not independent and can be obtained by combining these two.
We have a set of special solutions
$d_{1}=\frac{A}{2N},\quad d_{2}=B,\quad c_{4}=\frac{8NB^{2}}{A},$ (85)
where $A$ and $B$ are constants. As we will see, this set of special solutions will simplify the calculations in the subsequent discussion.
When we discuss the degeneracy of the matrix $\mathcal{B}_{(0)}^{ab}$, we need to be careful to identify which parts are weakly equal to zero. To clarify this, we first discuss how secondary constraints participate in the weak equality.
Since in the Lagrangian 58 the velocity term $\dot{h}_{ij}$ (encoded in $K_{ij}$) has been linearized, each canonical momentum corresponds to a primary constraint. Then all the secondary constraints can be combined into forms that are independent of canonical momenta. For example, we make a linear combination of the constraints $\tilde{\pi}^{ij}$ and $\mathcal{C}$,
$\mathcal{C}_{1}=2\tilde{\pi}^{ij}B_{ij}-\mathcal{C}=\frac{\delta S}{\delta N}-\frac{B_{ij}}{N}\frac{\delta S}{\delta B_{ij}}.$ (86)
By combining constraints $\mathcal{C}_{1}$ and $\frac{\delta S}{\delta\phi}$, we can obtain a constraint $\mathcal{C}_{2}$,
$\mathcal{C}_{2}$ $=$ $d_{2}\frac{\delta S}{\delta\phi}-2d_{1}N\mathcal{C}_{1}$ (87)
$=$ $N\sqrt{h}\left[\left(B\frac{\partial c_{1}}{\partial\phi}-A\Big{(}-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\Big{)}\right)B_{ij}B^{ij}\right.$
$\left.+\left(B\frac{\partial c_{2}}{\partial\phi}-A\Big{(}-\frac{c_{2}}{N}+\frac{\partial c_{2}}{\partial N}\Big{)}\right)B^{2}+\left(B\frac{\partial c_{3}}{\partial\phi}-A\Big{(}\frac{c_{3}}{N}+\frac{\partial c_{3}}{\partial N}\Big{)}\right)R\right],$
which does not depend on the special solution (85) chosen before.
Since we have replaced $\mathcal{B}^{21}$ with a term independent of the canonical momentum, (38) can be equivalently written as
$\det\mathcal{B}^{ab}(\vec{x},\vec{y})=\alpha(\vec{x},\vec{y})\mathcal{C}_{2}(\vec{x})+\beta(\vec{x},\vec{y})\frac{\delta S}{\delta\phi(\vec{x})}.$ (88)
By solving equations
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})\approx 0^{b},\qquad b=3,4,...,8,$ (89)
we obtain
$\left(\mathcal{V}_{1}\right)_{ij}=-2d_{1}N\mathcal{V}_{ij}^{(N)}+d_{2}\mathcal{V}_{ij}^{(\phi)},$ (90)
where
$\mathcal{V}_{ij}^{(N)}$ $=$ $\frac{1}{c_{1}}\frac{\partial c_{1}}{\partial N}B_{ij}+\left(\frac{1}{c_{1}+3c_{2}}\frac{\partial c_{2}}{\partial N}-\frac{c_{2}}{(c_{1}+3c_{2})c_{1}}\frac{\partial c_{1}}{\partial N}\right)B\delta_{ij},$ (91)
$\mathcal{V}_{ij}^{(\phi)}$ $=$ $\frac{1}{c_{1}}\frac{\partial c_{1}}{\partial\phi}B_{ij}+\left(\frac{1}{c_{1}+3c_{2}}\frac{\partial c_{2}}{\partial\phi}-\frac{c_{2}}{(c_{1}+3c_{2})c_{1}}\frac{\partial c_{1}}{\partial\phi}\right)B\delta_{ij}.$ (92)
Since we have previously chosen a set of special solutions for $d_{1}$, $d_{2}$, and $c_{4}$, equation (72) with $i=0$ can be written as
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})=D_{1}^{b}(\vec{y})B_{ij}(\vec{y})B^{ij}(\vec{y})+D_{2}^{b}(\vec{y})B^{2}(\vec{y})+D_{3}^{b}(\vec{y})R(\vec{y}),\qquad b=1,2.$ (93)
Since there are no terms such as $a_{i}a^{i}$ in equation (93), the coefficient $\beta(\vec{x},\vec{y})$ is fixed to zero. The above formula can equivalently be written as
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})=\alpha^{b}(\vec{y})\mathcal{C}_{2}(\vec{y}),\qquad b=1,2.$ (94)
In order for (94) to have a solution, the coefficients of $B_{ij}B^{ij}$, $B^{2}$, and $R$ must vanish. There will be 6 equations for $c_{1}$, $c_{2}$, and $c_{3}$. We assume that
$\frac{c_{1}}{c_{2}}=\mathrm{const}.$ (95)
This assumption simplifies the expression of $D_{2}^{b}$.
By fixing the coefficients of $B_{ij}B^{ij}$ and $R$ in (94), we obtain 4 differential equations
$A\left[\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{1}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 , (96)
$A\left[\frac{2}{N}\frac{\partial c_{1}}{\partial N}+\frac{\partial^{2}c_{1}}{\partial N^{2}}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial N}-\alpha^{2}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{2}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 ,
$A\left[\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{1}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0,0 ,
$A\left[\frac{2}{N}\frac{\partial c_{3}}{\partial N}+\frac{\partial^{2}c_{3}}{\partial N^{2}}-\alpha^{2}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{2}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0.0 . (97)
As long as equations (95) and (96)-(97) are satisfied, the coefficient of $B^{2}$ will be automatically zero. Therefore, we do not have to discuss the differential equation corresponding to the coefficient of $B^{2}$ here.
We also choose a set of special solutions
$c_{1}$ $=$ $Nf(\phi),$ (98)
$c_{2}$ $=$ $CNf(\phi),$ (99)
$c_{3}$ $=$ $\frac{D}{N}g(\phi),$ (100)
where the coefficients $C$ and $D$ are both constants.
By plugging the expressions of $c_{1}$ and $c_{3}$ into (96)-(97), we can simplify the four differential equations as
$\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (101)
$\alpha^{2}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (102)
$\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0,0 , (103)
$\alpha^{2}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0.0 . (104)
Finally, constraint $\mathcal{C}_{2}$ becomes
$\mathcal{C}_{2}$ $=BN\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B_{ij}B^{ij}+\frac{\partial c_{2}}{\partial\phi}B^{2}+\frac{\partial c_{3}}{\partial\phi}R\right).$ (105)
As we can see, if both $c_{1}$ and $c_{3}$ are independent of $\phi$, constraint $\mathcal{C}_{2}$ will not exist. The number of second-class constraints will be reduced by one, and the number of degrees of freedom will increase by 1/2. This is the linear-constraints branch of the 2DOF condition mentioned in the previous section, which is not the case we are interested in. Therefore, we assume that $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously. As a result, the coefficient $\alpha^{2}$ must be fixed to zero.
V.3 The second condition
According to the above discussion, after imposing the first 2DOF condition $\mathcal{S}_{1}$, we can obtain the constraint $\mathcal{C}^{\prime}$ from the null eigenvector $\mathcal{V}_{a}(\vec{x})$,
$\int\mathrm{d}^{3}x\,\mathcal{C}^{\prime}(\vec{x})V(\vec{x})\equiv\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}-2N\tilde{\pi}^{ij}&-\mathcal{C}&\frac{\delta S}{\delta\phi}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x}),$ (106)
which yields
$\mathcal{C}^{\prime}=BN\left[\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B^{ij}B_{ij}+\Big{(}\frac{2c_{2}}{c_{1}}\frac{\partial c_{1}}{\partial\phi}-\frac{\partial c_{2}}{\partial\phi}\Big{)}B^{2}-\frac{\partial c_{3}}{\partial\phi}R\right)-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial\phi}\pi^{ij}B_{ij}\right].$ (107)
The Poisson brackets between the constraint $\mathcal{C}^{\prime}$ and the constraints $p$, $\pi$, and $p^{kl}$ are
$\left[\mathcal{C}^{\prime}(\vec{x}),p(\vec{y})\right]$ $=\left[-BN\frac{\partial}{\partial\phi}\left(\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial\phi}\right)\tilde{\pi}^{ij}B_{ij}-\frac{\partial c_{3}}{\partial\phi\partial\phi}\left(\frac{\partial c_{3}}{\partial\phi}\right)^{-1}\mathcal{C}_{2}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (108)
$\left[\mathcal{C}^{\prime}(\vec{x}),\pi(\vec{y})\right]=\frac{B}{N(\vec{x})}\left(\mathcal{C}^{\prime}(\vec{x})+\mathcal{C}_{2}(\vec{x})\right)\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (109)
$\left[\mathcal{C}^{\prime}(\vec{x}),p^{kl}(\vec{y})\right]$ $=-\frac{BN(\vec{x})}{c_{1}(\vec{x})}\frac{\partial c_{1}(\vec{x})}{\partial\phi(\vec{x})}\tilde{\pi}^{kl}(\vec{x})\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (110)
which all vanish on the constraint surface. The Poisson bracket between the constraint $\mathcal{C}^{\prime}$ and itself is generally nonzero, unless
$\frac{\partial c_{1}}{\partial\phi}=0,\quad\text{or}\quad\frac{\partial c_{3}}{\partial\phi}=0.$ (111)
As mentioned above, $c_{1}$ and $c_{3}$ cannot both be independent of $\phi$ simultaneously; otherwise, the number of degrees of freedom will increase by 1/21/21 / 2. If we require the second 2DOF condition $\mathcal{S}_{2}^{(2)}$ to hold, there are two cases.
In the first case, if we choose $\frac{\partial c_{1}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
$\mathcal{C}_{2}=BN\sqrt{h}\frac{\partial c_{3}}{\partial\phi}R\approx 0.$ (112)
This case is clearly non-physical, as it fixes $R$ to vanish.
On the other hand, if we choose $\frac{\partial c_{3}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
$\mathcal{C}_{2}$ $=BN\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B_{ij}B^{ij}+\frac{\partial c_{2}}{\partial\phi}B^{2}\right)\approx 0.$ (113)
In this case, we can solve for all the coefficients:
$c_{1}$ $=Nf(\phi),\quad c_{2}=CNf(\phi),\quad c_{3}=\frac{D}{N},$
$c_{4}$ $=\frac{8NB^{2}}{A},\quad d_{1}=\frac{A}{2N},\quad d_{2}=B.$ (114)
The resulting theory thus corresponds to a SCG with an auxiliary scalar field of $d=2$ that propagates two degrees of freedom. The constraint coefficients $\alpha^{i}$ in the previous differential equation are also uniquely determined:
$\alpha^{1}=-\left(\frac{\partial f(\phi)}{\partial\phi}\right)^{-1}\left(\frac{2}{f(\phi)}\Big{(}\frac{\partial f(\phi)}{\partial\phi}\Big{)}^{2}-\frac{\partial^{2}f(\phi)}{\partial\phi^{2}}\right),\quad\alpha^{2}=0.$ (115)
Now let us consider the sub-matrix $\mathcal{E}^{ab}$ of the simplified Dirac matrix $\mathcal{M}^{\prime}{}^{ab}$,
$\mathcal{E}^{ab}(\vec{x},\vec{y})=\begin{array}[]{|c|cccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\mathcal{C}^{\prime}(\vec{y})\\ \hline\cr\mathcal{C}^{\prime}(\vec{x})&0&0&0&0\\ \tilde{\pi}^{ij}(\vec{x})&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X&X\\ \Phi(\vec{x})&X&X&X&X\\ \hline\cr\end{array},$ (116)
whose determinant is obviously zero. Based on the discussion of null eigenvectors in Appendix C, we can find a null eigenvector $\mathcal{V}_{a}^{(\mathcal{E})}$ to construct a new first-class constraint $\tilde{\mathcal{C}}$,
$\int\mathrm{d}^{3}x\,\tilde{\mathcal{C}}(\vec{x})V(\vec{x})=\int\mathrm{d}^{3}x\left(\begin{array}[]{cccc}p&\pi&p^{kl}&\mathcal{C}^{\prime}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}{}_{1}\right)_{ij}^{(\mathcal{E})}&\mathcal{\mathcal{V}}{}_{2}^{(\mathcal{E})}&\mathcal{\mathcal{V}}{}_{3}^{(\mathcal{E})}\end{array}\right)(\vec{x})\approx 0.$ (117)
To summarize, there are 24 constraints which can be divided into two classes:
$\begin{array}[]{rl}7\text{ first-class: }&\pi_{i},\quad\mathcal{C}_{i},\quad\tilde{\mathcal{C}},\\ 16\text{ second-class: }&p,\quad p^{ij},\quad\pi,\quad\tilde{\pi}^{ij},\quad\frac{\delta S}{\delta\phi},\quad\Phi.\end{array}$
The number of DOFs is thus
#$\#_{\text{dof}}$ $=\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (118)
[…]
=2=2.= 2 .
Finally, let us discuss another 2DOF condition $\mathcal{S}_{2}^{(1)}$. By applying the consistency condition of the constraint $\mathcal{C}_{2}$, we can obtain the constraint $\Phi$,
$\Phi(\vec{x})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),H_{C}\right]$ (119)
$=$ $-B\frac{2N(\vec{x})}{c_{1}(\vec{x})}\frac{\partial c_{1}(\vec{x})}{\partial\phi(\vec{x})}B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}$
$-\int d^{3}z\,BN(\vec{z})\sqrt{h}\frac{\delta}{\delta h_{ij}(\vec{z})}\left(\frac{\partial c_{1}}{\partial\phi}B^{ij}B_{ij}+\Big{(}\frac{2c_{2}}{c_{1}}\frac{\partial c_{1}}{\partial\phi}-\frac{\partial c_{2}}{\partial\phi}\Big{)}B^{2}-\frac{\partial c_{3}}{\partial\phi}R\right)(\vec{x})2N(\vec{z})B_{ij}(\vec{z}).$
Since the second-order derivative of the delta function arises when calculating the integral in the last line, the constraint $\Phi$ contains $\mathrm{D}^{i}\mathrm{D}_{j}B_{kl}$, which is proportional to
$\frac{\partial c_{3}}{\partial\phi}\mathrm{D}^{i}\mathrm{D}_{j}B_{kl}.$ (120)
Therefore, requiring the 2DOF condition $\mathcal{S}_{2}^{(1)}=\det\mathcal{D}^{ab}\approx 0$ also implies that $\frac{\partial c_{3}}{\partial\phi}$ may vanish as one of the conditions for the matrix $\mathcal{D}^{ab}$ to be degenerate, which is sufficient to satisfy the 2DOF condition $\mathcal{S}_{2}^{(2)}$.
VI Conclusion
In this work, we investigated a class of spatially covariant gravity theories with a non-dynamical scalar field. We briefly described our model in Sec. II. The general Lagrangian is given by (3), which has been shown to have two tensorial and one scalar DOFs [85]. The purpose of this work is to determine the conditions on the Lagrangian, which we dub the 2DOF conditions, under which only two degrees of freedom propagate.
A perturbative approach was taken in [86] to derive conditions such that the scalar mode is eliminated at linear order in perturbations. In this work, we employ a strict Hamiltonian constraint analysis to derive conditions such that only two DOFs propagate in the nonperturbative sense. In Sec. III, we describe the Hamiltonian formalism for the primary constraints and their consistency conditions. As expected, there are 3 DOFs in the theory if no further conditions are imposed.
In Sec. IV, by requiring the degeneracy of the Dirac matrix, we find that two conditions are required to make the theory propagate only 2 DOFs. The first condition (38) states that the matrix $\mathcal{B}^{ab}$ given in (39) must be degenerate, which results in a secondary constraint. To fully eliminate the scalar mode, a second condition is necessary. The second condition can be divided into two categories according to their different effects on the Dirac matrix. In the first case, the condition (49) generally turns a second-class constraint into a first-class constraint. In the second case, the condition (50) generates another secondary constraint $\Phi$.
We would like to emphasize that even if the Lagrangian satisfies both 2DOF conditions, the number of DOFs does not necessarily decrease. A special case is when the constraints are linearly dependent, which implies that the number of DOFs may remain unchanged. Another special case, which we refer to as the non-physical branch of the 2DOF conditions, renders the theory trivial or physically unacceptable (e.g., lacking tensor modes). One must therefore be careful when dealing with the 2DOF conditions in order to select the physical case, i.e., theories with precisely two degrees of freedom.
In Sec. V, based on the spatially covariant monomials classified in [57], [58], we consider a concrete Lagrangian (60) built from monomials of $d=2$ (i.e., with two derivatives) as an illustration of our formal analysis. Since the Dirac matrix of this model contains derivatives of the delta function, it is complicated to calculate the matrix form of the 2DOF conditions for this concrete model. Instead, we choose an equivalent but more efficient method to obtain the 2DOF conditions, which is to find the null eigenvector $\mathcal{V}_{a}$ satisfying (40). The resulting 2DOF conditions are given in (66) and (111). In particular, for this concrete model, we are able to eliminate the non-physical branch and the linear constraints branch of the 2DOF conditions. Finally, we find a set of coefficients (114), and the resulting Lagrangian corresponds to a spatially covariant gravity theory with an auxiliary scalar field that propagates only two DOFs.
Acknowledgements.
We would like to thank Zhi-Chao Wang for valuable discussions. X.G. is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12475068 and No. 11975020 and the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2025A1515012977.
Appendix A Comparison with scalar-tensor theory in the spatial gauge
The spatially covariant gravity with an auxiliary scalar field was firstly proposed in [85], where the Hamiltonian analysis was performed to show that the theory propagates 3 degrees of freedom. This idea was originally motivated by generally covariant scalar-tensor theory when the scalar field possesses a spacelike gradient. After choosing the so-called “spatial gauge” as in [85], the resulting action takes a form similar to (1). However, we should note that “spatially covariant gravity with an auxiliary scalar field” and “scalar-tensor theory in the spatial gauge” are completely different theories.
First, we will show the difference between (1) and the scalar-tensor theory in the spatial gauge. Let us consider a general action of scalar-tensor theory
$S_{\mathrm{GST}}=\int\mathrm{d}^{4}x\sqrt{-g}\mathcal{L}\left(\phi;g_{ab},\varepsilon_{abcd},{}^{4}\!R_{abcd};\nabla_{a}\right),$ (121)
where the Lagrangian involves a scalar field $\phi$, the spacetime metric $g_{ab}$, the spacetime curvature tensor ${}^{4}\!R_{abcd}$ as well as their covariant derivatives. The 4-dimensional Levi-Civita tensor $\varepsilon_{abcd}$ encodes possible parity violation effects.
In a cosmological context, the scalar field $\phi$ is assumed to possess a timelike gradient so that the so-called unitary gauge with $\phi=t$ can be chosen. Here we consider the opposite situation by assuming that the scalar field possesses a spacelike gradient [^3] Similar ideas appear in “elastic inflation” [87] and “solid inflation” [88], where it is the Nambu-Goldstone boson breaking spatial diffeomorphism that plays the role of the scalar field.. Contrary to the unitary gauge which is defined by requiring $\mathrm{D}_{a}\phi=0$, we now choose a gauge in which
£$\pounds_{\bm{n}}\phi=0,$ (122)
where $n^{a}$ is the normal vector to the hypersurfaces. This can be understood as choosing hypersurfaces such that the normal vector $n^{a}$ exactly lies on the constant $\phi$ hypersurfaces. As a result, the value of $\phi$ does not change when transported along the normal vector. This choice of spatial hypersurfaces is referred to as the “spatial gauge” in [85].
The normal vector $n^{a}$ is normalized by $n^{a}n_{a}=-1$. The induced metric on the spatial hypersurfaces is, as usual,
$h_{ab}=g_{ab}+n_{a}n_{b}.$ (123)
We can then split all the 4-dimensional objects into their temporal and spatial parts. For example, the decomposition of $\nabla_{a}\phi$ is
[…]
where $\mathrm{D}_{a}$ is the covariant derivative compatible with the induced metric $h_{ab}$. In the last equality of (124) we have used the fact that £$\pounds_{\bm{n}}\phi=0$ in the spatial gauge. This is exactly the opposite of the unitary gauge, where $\mathrm{D}_{a}\phi=0$ and thus $\nabla_{a}\phi\rightarrow-n_{a}\pounds_{\bm{n}}\phi$.
After performing the 3+1 decomposition and choosing the spatial gauge, the action (121) can be written in the form
$S_{\mathrm{GST}}^{\mathrm{(s.g.)}}=\int\mathrm{d}t\mathrm{d}^{3}x\,N\sqrt{h}\,\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij};\mathrm{D}_{i},\pounds_{\bm{n}}\right),$ (125)
where “s.g.” stands for the spatial gauge and we have fixed the spatial coordinates adapted to the spacelike hypersurfaces. At first glance, the action (125) takes the same form as (1); however, the crucial difference is that in the spatial gauge, the scalar field $\phi$ can be viewed as a time-independent but space-dependent field, which breaks spatial diffeomorphism. In other words, $\phi=\phi(\bm{x})$ in (125) cannot be viewed as a dynamical nor auxiliary variable, but is actually a “function” of spatial coordinates with fixed values. This is completely different from (1), which has spatial covariance and where $\phi$ plays the role of an auxiliary field.
Next, we show that the generally covariant version (i.e., the correspondence) of (1) is nothing but a “two-field” scalar-tensor theory. Let us take $\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi$ as an example. The covariant version is
$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi=h^{ij}\mathrm{D}_{i}\phi\mathrm{D}_{j}\phi\rightarrow\left(g^{ab}+u^{a}u^{b}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (126)
where $u_{a}=-N\nabla_{a}\Phi$ with $\Phi$ being the scalar field defining the spacelike hypersurfaces. In other words, the spacelike hypersurfaces are defined by $\Phi=\mathrm{const.}$. In (126), $N=1/\sqrt{-\left(\nabla\Phi\right)^{2}}$, which reduces to the lapse function when fixing the so-called unitary gauge with $\Phi=t$. By expanding (126) explicitly, we get
$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi\rightarrow\left(g^{ab}-\frac{\nabla^{a}\Phi\nabla^{b}\Phi}{\left(\nabla\Phi\right)^{2}}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (127)
which is clearly a two-field scalar-tensor theory term.
Appendix B Constraint algebra
In this appendix, we present the explicit expressions for the Poisson brackets.
The Poisson brackets among constraints are
$\left[p(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ $=$ $\frac{1}{2N(\vec{y})}\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta B_{kl}(\vec{y})},$ (128)
$\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ $=$ $\frac{1}{2}\frac{1}{N(\vec{y})}\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})},$ (129)
$\left[\pi(\vec{x}),\tilde{\pi}^{ij}(\vec{y})\right]$ $=$ $-\frac{1}{2}\delta^{3}(\vec{x}-\vec{y})\frac{1}{N^{2}(\vec{y})}\frac{\delta S}{\delta B_{ij}(\vec{y})}+\frac{1}{2}\frac{1}{N(\vec{y})}\frac{\delta^{2}S}{\delta N(\vec{x})\delta B_{ij}(\vec{y})},$ (130)
$\left[\tilde{\pi}^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ $=$ $\frac{1}{2N(\vec{y})}\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta B_{kl}(\vec{y})}-\frac{1}{2N(\vec{x})}\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{y})},$ (131)
$[\pi(\vec{x}),\mathcal{C}(\vec{y})]$ $=$ $\frac{\delta^{2}S}{\delta N(\vec{x})\delta N(\vec{y})},$ (132)
${}[p(\vec{x}),\mathcal{C}(\vec{y})]$ $=$ $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta N(\vec{y})},$ (133)
$\left[p^{ij}(\vec{x}),\mathcal{C}(\vec{y})\right]$ $=$ $-2\delta^{3}(\vec{x}-\vec{y})\pi^{ij}(\vec{y})+\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta N(\vec{y})},$ (134)
$\left[\tilde{\pi}^{ij}(\vec{x}),\mathcal{C}(\vec{y})\right]$ $=$ $\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta N(\vec{y})}-\frac{1}{N(\vec{x})}\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{y})}B_{kl}(\vec{y}),$ (135)
$\left[p(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})},$ (136)
$\left[p^{ij}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})},$ (137)
$\left[\pi(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta N(\vec{x})\delta\phi(\vec{y})},$ (138)
$\left[\tilde{\pi}^{ij}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta\phi(\vec{y})},$ (139)
$[\mathcal{C}(\vec{x}),\mathcal{C}(\vec{y})]$ $=$ $-\frac{\delta^{2}S}{\delta h_{ij}(\vec{y})\delta N(\vec{x})}2B_{ij}(\vec{y})+2B_{ij}(\vec{x})\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta N(\vec{y})},$ (140)
$\left[\frac{\delta S}{\delta\phi(\vec{x})},\mathcal{C}(\vec{y})\right]$ $=$ $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta h_{ij}(\vec{y})}2B_{ij}(\vec{y}).$ (141)
The Poisson brackets involving the canonical Hamiltonian are
$\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $\frac{\delta S}{\delta h_{ij}(\vec{x})}-\frac{1}{N(\vec{x})}\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{z})}N(\vec{z})B_{kl}(\vec{z}),$ (142)
$\left[\mathcal{C}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $2B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}-\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta N(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}),$ (143)
$\left[\frac{\delta S}{\delta\phi(\vec{x})},H_{\mathrm{C}}\right]$ $=$ $\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}).$ (144)
Appendix C Adjoint matrix form of null eigenvector
If we wish to find the null eigenvector of the degenerate matrix $\mathcal{B}^{ab}$, one method is to calculate the non-zero rows (or columns) of its corresponding adjoint matrix $\mathcal{B}_{ab}^{*}$, and to multiply the corresponding elements by $(-1)^{a+b}$ to build the null eigenvector. For simplicity, we assume that $\mathcal{B}^{ab}$ is a simple $\times 3$ 3 matrix and the rows and columns correspond to constraints $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$.
To verify that this form of null eigenvector is feasible, we assume that the matrix $\mathcal{B}^{ab}$ is degenerate and without loss of generality we take the first row of its adjoint matrix $\mathcal{B}_{ab}^{*}$ to form a null eigenvector
$\mathcal{V}_{a}=\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right).$ (145)
We then multiply the null eigenvector $\mathcal{V}_{a}$ by the matrix $\mathcal{B}_{ab}^{*}$,
$\mathcal{V}_{a}\mathcal{B}^{ab}=\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{1b}&\mathcal{B}^{2b}&\mathcal{B}^{3b}\end{array}\right)^{T},$ (146)
which yields
$\mathcal{V}_{a}\mathcal{B}^{a1}$ $=$ $\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{11}&\mathcal{B}^{21}&\mathcal{B}^{31}\end{array}\right)^{T}$ (149)
$=$ $\det\mathcal{B}^{ab}=0,$ (150)
$\mathcal{V}_{a}\mathcal{B}^{a2}$ $=$ $\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{12}&\mathcal{B}^{22}&\mathcal{B}^{32}\end{array}\right)^{T}$ (153)
$=$ $\det\left(\begin{array}[]{ccc}\mathcal{B}^{12}&\mathcal{B}^{12}&\mathcal{B}^{13}\\ \mathcal{B}^{22}&\mathcal{B}^{22}&\mathcal{B}^{23}\\ \mathcal{B}^{32}&\mathcal{B}^{32}&\mathcal{B}^{33}\end{array}\right)=0,$ (157)
$\mathcal{V}_{a}\mathcal{B}^{a3}$ $=\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{13}&\mathcal{B}^{23}&\mathcal{B}^{33}\end{array}\right)^{T}$ (160)
$=\det\left(\begin{array}[]{ccc}\mathcal{B}^{13}&\mathcal{B}^{12}&\mathcal{B}^{13}\\ \mathcal{B}^{23}&\mathcal{B}^{22}&\mathcal{B}^{23}\\ \mathcal{B}^{33}&\mathcal{B}^{32}&\mathcal{B}^{33}\end{array}\right)=0.$ (164)
We can combine constraints $\phi_{i}$ with the null eigenvectors to make a new constraint
$\phi^{\prime}(\vec{y})=\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}\phi_{1}&\phi_{2}&\phi_{3}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)^{T}(\vec{x},\vec{y}).$ (165)
The Poisson brackets of $\phi^{\prime}$ with constraints $\phi_{i}$ are all vanishing,
$\left[\phi_{i}(\vec{x}),\phi^{\prime}(\vec{y})\right]$$\approx$ $\int\mathrm{d}^{3}z\left[\phi_{i}(\vec{x}),\left(\begin{array}[]{ccc}\phi_{1}&\phi_{2}&\phi_{3}\end{array}\right)(\vec{z})\right]\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)^{T}(\vec{z},\vec{y})$ (168)
$=$ $\int\mathrm{d}^{3}z\left(\begin{array}[]{ccc}\mathcal{B}^{1i}&\mathcal{B}^{2i}&\mathcal{B}^{3i}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)^{T}(\vec{z},\vec{y})$ (171)
$=$ 0 0.0 . (172)
In practical calculations, in order to avoid integrating the delta function of higher order, we need to multiply the null eigenvector by a greatest common divisor, i.e.,
$\mathcal{V}_{a}=\left(\begin{array}[]{ccc}\mathcal{\mathcal{V}}_{1}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)=\frac{1}{\mathrm{gcd}(\mathcal{B}_{1i}^{*},\mathcal{B}_{2i}^{*},\mathcal{B}_{3i}^{*})}\left(\begin{array}[]{ccc}\mathcal{B}_{1i}^{*}&-\mathcal{B}_{2i}^{*}&\mathcal{B}_{3i}^{*}\end{array}\right).$ (173)

### v2

@@
 I Introduction
-With the detection of gravitational waves, a novel tool has emerged for testing theories of gravity.
-A fundamental question arises: Is general relativity the only theory of gravity that propagates gravitational waves with precisely two degrees of freedom (2DOFs)?
-This question is partially addressed by Lovelock’s theorem [1, 2], which states that in four dimensions, if one requires spacetime covariance and second-order equations of motion, general relativity is indeed the unique theory that propagates gravitational waves with two degrees of freedom.
-This also implies that, by violating the assumptions of Lovelock’s theorem, other gravitational theories propagating two degrees of freedom can, in principle, exist.
-Gravitational theories with 2DOFs can be traced back to the Cuscuton theory [3], which has been extensively studied [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], and extended in the framework of scalar-tensor theories with higher-order derivatives [17, 18].
-Such theories have also been discussed as special cases of Hořava gravity [19, 20, 21, 22].
-The connection between 2DOFs and spacetime symmetries has been investigated in [23, 24, 25, 26].
-Another class of theories propagating 2DOFs was proposed in [27], known as minimally modified gravity (MMG) [^1] See also the minimal theory of massive gravity studied in [28, 29, 30, 31, 32, 33]..
+In light of the detection of gravitational waves, gravitational waves provide a novel tool for testing theories of gravity.
+General relativity possesses only two propagating degrees of freedom (DOFs), which manifests as tensorial gravitational waves (GWs) in a homogeneous and isotropic background.
+The detection of GWs raises a fundamental question: Is general relativity the only theory of gravity that propagates GWs with precisely two degrees of freedom (2DOFs)?
+This question is partially answered by Lovelock’s theorem [1, 2], which states that in four dimensions, requiring spacetime covariance and second-order equations of motion, general relativity is indeed the unique theory propagating gravitational waves with two degrees of freedom.
+This also implies that, by violating the assumptions of Lovelock’s theorem, other theories of gravity propagating 2DOFs can, in principle, exist.
+From the theoretical point of view, a Lorentz covariant massless spin-2 particle possesses two physical polarization states.
+If one constructs a Lorentz invariant theory for such a massless spin-2 particle and further demands that it can be consistently coupled to matter fields, the resulting structure is necessarily that of general relativity.
+In this sense, general relativity as a theory of gravity is unique since its form is inextricably related to the masslessness of the graviton and the Lorentz symmetry of spacetime.
+Consequently, alternative theories of gravity that also propagate only 2DOFs provide a valuable testing ground for the foundation principles of general relativity, such as Lorentz invariance, diffeomorphism invariance, and the masslessness of graviton, etc.
+Gravitational theories propagating 2DOFs can be traced back to the Cuscuton theory [3], which is introduced as a scalar field model with a non-canonical kinetic term $\propto\sqrt{\left|-\partial_{\mu}\phi\partial^{\mu}\phi\right|}$.
+The cuscuton theory represents the incompressible (infinite speed of sound) limit of $k$-essence, in which the scalar perturbation propagates with an infinite speed of sound in a cosmological background.
+Although linear perturbations exhibit superluminal propagation, the local phase space degenerates, implying no independent dynamical DOFs and thus preserving causality.
+As a result, the cuscuton theory modifies gravitational dynamics without introducing additional local DOFs.
+The Cuscuton theory was extensively studied [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and was extended in the framework of scalar-tensor theory with higher order derivatives [17, 18].
+Such kind of theory has also been discussed as a special case of Hořava gravity [19, 20, 21, 22].
+The relation between the 2DOFs and the spacetime symmetry has been investigated in [23, 24, 25, 26].
+Another class of of theories propagating 2DOFs was proposed in [27], which is dubbed the minimally modified gravity (MMG) [^1] See also the minimal theory of massive gravity studied in [28, 29, 30, 31, 32, 33]..
+As the name suggests, MMG theories modify generally relativity “minimally” in the sense that they modify the gravitational sector without introducing extra DOFs beyond the standard 2DOFs.
+MMG theories extend general relativity by breaking time diffeomorphism invariance while preserving spatial diffeomorphism invariance, and whose action is linear in the lapse function.
+A key result is the derivation of a self-consistency condition, which ensures the theory possesses no more than two local physical DOFs.
+Examples of MMG include general relativity, square-root gravity, exponential gravity, and theories with lapse-independent terms, all of which exhibit a constrained phase space structure that eliminates the scalar graviton through first- or second-class constraints.
 The so-called type-II MMG [34, 35] includes the original Cuscuton theory as a special case.
-MMG has been extensively studied [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], and has also been generalized by introducing auxiliary constraints in phase space [52, 53].
-A general framework for spatially covariant gravity (SCG), which respects only spatial covariance, was proposed in [54, 55].
-Due to the breaking of full spacetime covariance, SCG generally propagates a scalar degree of freedom in addition to the 2DOFs corresponding to gravitational waves.
-Thus, it provides an alternative approach to constructing scalar-tensor theories [56, 57, 58, 59].
-SCG includes Hořava gravity [60, 61, 62] and the effective field theory of inflation/dark energy [63, 64, 65, 66, 67] as special cases.
-It has been further generalized to include a dynamical lapse function [68, 69], nonmetricity [70], parity violation [71], and multiple scalar modes [72].
-SCG has also been applied in cosmology and gravitational wave studies [73, 74, 75, 76, 77].
-Subclasses of SCG that propagate only 2DOFs were explored in [78, 79], and in the presence of a dynamical lapse function in [80].
-The explicit Lagrangian found in [78] has been employed in cosmological studies [81, 82, 83, 10, 84].
-In this work, we generalize the SCG framework to include an auxiliary scalar field and determine the conditions under which the theory propagates only 2DOFs.
-The idea of spatially covariant gravity with an auxiliary scalar field was first introduced in [85], originally motivated by generally covariant scalar-tensor theories in which the scalar field has a spacelike gradient [^2] However, we emphasize that these are entirely different theories.
-See Appendix A for details..
-The presence of an auxiliary scalar field introduces new and novel possibilities for constructing 2DOF theories.
-In [86], a perturbative analysis similar to [17, 69, 79] was performed to derive necessary conditions for the absence of scalar modes at linear order in perturbations around a cosmological background.
-Nevertheless, a Hamiltonian constraint analysis is still required to determine the necessary and sufficient conditions for the theory to propagate only 2DOFs at the nonlinear level.
-This paper is devoted to this task.
-The paper is organized as follows.
+The MMG has been studied extensively [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51].
+It has also been generalized in [52, 53] by introducing auxiliary constraints in the phase space.
+A general framework of spatially covariant gravity (SCG) theories respecting only the spatial covariance was proposed in [54, 55].
+Due to the violation of general covariance, SCG generally propagates a scalar degree of freedom (besides the 2DOFs corresponding to the gravitational waves) and thus can be viewed as an alternative approach to constructing the scalar-tensor theories [56, 57, 58, 59].
+The SCG includes the Horava gravity [60, 61, 62], effective field theory of inflation/dark energy [63, 64, 65, 66, 67] as special cases.
+It was further generalized with a dynamical lapse function [68, 69], with nonmetricity [70], with parity violation [71] and with multiple scalar modes [72].
+The SCG has been applied in the study of cosmology and gravitational waves [73, 74, 75, 76, 77].
+Subclasses of SCG propagating only 2DOFs were explored in [78, 79] and with a dynamical lapse function in [80].
+The concrete Lagrangian found in [78] has been applied in the study of cosmology [81, 82, 83, 10, 84].
+In this work, we will generalize the SCG in the presence of an auxiliary scalar field, and determine the conditions such that the theory propagates only 2DOFs.
+The idea of spatially covariant gravity with an auxiliary scalar field was firstly introduced in [85], which was originally motivated by the generally covariant scalar-tensor theory when the scalar field possess a spacelike gradient [^2] However, we would like to point out that they are completely different theories.
+Please refer to Appendix A for details.
+.
+The presence of an auxiliary scalar field gives us more and novel possibilities to build the theory propagating two degrees of freedom.
+In [86], by performing a perturbative analysis similar to [17, 69, 79], necessary conditions such that no scalar mode propagates at linear order in perturbations around a cosmological background were determined.
+Nevertheless a Hamiltonian constraint analysis is still needed in order to determine the necessary and sufficient conditions for the theory to propagate only 2DOFs at the nonlinear level.
+This paper is thus devoted to this issue.
+This paper is organized as follows.
 In Sec.
 II, we describe our model of spatially covariant gravity with an auxiliary scalar field.
-We find that new terms in the Lagrangian are allowed due to the inclusion of the auxiliary scalar.
+It was found that new terms in the Lagrangian are allowed thanks to the introduction of the auxiliary scalar field.
 In Sec.
-III, we derive the Hamiltonian formalism and show that, in general, a scalar mode propagates unless additional constraints are imposed.
+III, we derive the Hamiltonian formalism and show that the theory generally propagates a scalar mode if no further constraint is imposed.
 In Sec.
-IV, we perform a degeneracy analysis and derive the conditions under which the theory propagates only 2DOFs.
+IV, we make the degeneracy analysis and derive the conditions such that the theory propagate only 2DOFs.
 In Sec.
-V, we apply our formalism to a concrete model, where the Lagrangian is built from SCG monomials with $d=2$.
-Finally, in Sec.
+V, we apply our formal analysis to a concrete model, of which the Lagrangian is built of SCG monomials of $d=2$.
+In Sec.
 VI, we summarize our results.
@@
 $S=\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij};\mathrm{D}_{i},\pounds_{\bm{n}}\right),$ (1)
-where $N$ is the usual lapse function, $h_{ij}$ is the spatial metric on the spacelike hypersurfaces, ${}^{3}\!R_{ij}$ is the spatial Ricci tensor, $\mathrm{D}_{i}$ is the spatially covariant derivative compatible with $h_{ij}$, and £$\pounds_{\bm{n}}$ is the Lie derivative with respect to the normal vector $n^{a}$ of the hypersurfaces.
-By definition, the action (1) respects only the three-dimensional spatial invariance.
-We require that the Lie derivative £$\pounds_{\bm{n}}$ acts only on $h_{ij}$, which yields the extrinsic curvature $K_{ij}$ defined by
-£$\pounds_{\bm{n}}h_{ij}=2K_{ij}=\frac{1}{N}\left(\dot{h}_{ij}-\pounds_{\vec{N}}h_{ij}\right).$ (2)
-As a result, both the lapse function $N$ and the scalar field $\phi$ play the role of auxiliary variables.
-Note that the shift vector $N^{i}$ should not appear explicitly in the Lagrangian, as it is actually the gauge field of spatial covariance.
-Since the Lie derivative enters the Lagrangian only through the extrinsic curvature, we simply introduce $K_{ij}$ to replace the Lie derivative £$\pounds_{\bm{n}}$ in (1), and consider the following action:
+where $N$ is the usual lapse function, $h_{ij}$ is the spatial metric on the spacelike hypersurfaces, $^{3}\!R_{ij}$ is the spatial Ricci tensor, $\mathrm{D}_{i}$ is the spatially covariant derivative compatible with $h_{ij}$, and $\pounds_{\bm{n}}$ is the Lie derivative with respect to the normal vector $n^{a}$ of the hypersurfaces.
+By definition, the action (1) respects only the 3-dimensional spatial invariance.
+We require that the Lie derivative $\pounds_{\bm{n}}$ acts only on $h_{ij}$, which yields the extrinsic curvature $K_{ij}$ defined by
+$\pounds_{\bm{n}}h_{ij}=2K_{ij}=\frac{1}{N}\left(\dot{h}_{ij}-\pounds_{\vec{N}}h_{ij}\right).$ (2)
+As a result, both the lapse function $N$ and the scalar field $\phi$ play the role of an auxiliary variables.
+Note the shift vector $N^{i}$ should not appear explicitly in the Lagrangian, which is actually the gauge field of the spatial covariance.
+In this work, we restrict our attention to the first-order Lie derivative to prevent higher-order time derivatives from appearing in the equations of motion.
+Moreover, since the Lie derivative is assumed to act solely on $h_{ij}$, it enters the Lagrangian exclusively through the extrinsic curvature (2).
+Therefore, we simply introduce $K_{ij}$ to replace the Lie derivative $\pounds_{\bm{n}}$ in (1), and consider the following action
 $\tilde{S}=\int\mathrm{d}t\mathrm{d}^{3}x\,N\sqrt{h}\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right).$ (3)
-According to the Hamiltonian analysis in the next section, the action (3) generally has 3 DOFs.
-This is simply because $\phi$ has no dynamics; the number of DOFs in (3) is the same as in spatially covariant gravity without the scalar field [54, 55].
-In principle, Lie derivatives of the lapse function £$\pounds_{\bm{n}}N$ and of the extrinsic curvature £$\pounds_{\bm{n}}K_{ij}$ can be considered, which would introduce additional degrees of freedom.
-It has been shown that extra conditions must be imposed in order to eliminate these unwanted degrees of freedom.
-For example, the theory will propagate an extra DOF (i.e., a total of 4 DOFs) in the case of a dynamical lapse function.
-Two conditions are required to eliminate such an extra DOF [68].
-A similar analysis was carried out to fully eliminate the scalar DOFs [80].
-Generally, the Lagrangian in (3) may be nonlinear in $K_{ij}$, which makes the explicit inversion of velocity $\dot{h}_{ij}$ in terms of the momentum $\pi^{ij}$ impossible.
-This problem can be solved by rewriting (3) in an equivalent form:
+According to the Hamiltonian analysis in the next section, generally the action (3) has 3 DOFs.
+This is simply because $\phi$ has no dynamics, the number of DOFs of (3) is the same as the spatially covariant gravity without the scalar field [54, 55].
+In particular, since the extrinsic curvature is linear in the Lie derivative (and thus in time derivative), the resulting equations of motion from the action (3) contain at most second-order time derivatives.
+As a result, the theory avoids the Ostrogradsky ghost instability.
+In principle, Lie derivatives of the lapse function $\pounds_{\bm{n}}N$ and the extrinsic curvature $\pounds_{\bm{n}}K_{ij}$ can be considered, which will introduced more degrees of freedom.
+It was shown that extra conditions must be put in order to evade these unwanted degrees of freedom.
+For example, the theory will propagate an extra DOF (i.e., in total 4 DOFs) in the case of a dynamical lapse function.
+Two conditions are needed in order to evade such an extra DOF [68].
+Similar analysis was considered in order to fully eliminate the scalar DOFs [80].
+Generally, the Lagrangian in (3) may be nonlinear in $K_{ij}$, which makes the explicit reversion of velocity $\dot{h}_{ij}$ in terms of the momentum $\pi^{ij}$ impossible.
+This problem can be solved by rewriting (3) in an equivalent form
 $\tilde{S}$ $=\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\tilde{\mathcal{L}}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right)$
 $=S\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\varLambda^{ij}\left(K_{ij}-B_{ij}\right),$ (4)
-where the action $S\left(\phi,N,h_{ij},{}^{3}R_{ij},B_{ij};\mathrm{D}_{i}\right)$ is the same as (3) with $K_{ij}$ replaced by $B_{ij}$.
-This strategy of using a Lagrange multiplier to linearize the velocities was employed in [68, 78].
-The equations of motion for $\varLambda^{ij}$ ensure that the auxiliary field $B_{ij}$ is exactly equal to $K_{ij}$.
-Thus, the two actions are equivalent, at least at the classical level.
-By making use of the equation of motion for the auxiliary field $B_{ij}$, we can determine the Lagrange multiplier $\varLambda^{ij}$ as
+where the action $S\left(\phi,N,h_{ij},{}^{3}R_{ij},B_{ij};\mathrm{D}_{i}\right)$ is the same action as (3) with $K_{ij}$ replaced by $B_{ij}$.
+This strategy of using the Lagrange multiplier to linearize the velocities has been used in [68, 78].
+The equations of motion for $\varLambda^{ij}$ ensure the auxiliary field $B_{ij}$ to be exactly identical to $K_{ij}$.
+Thus the two actions are equivalent, at least at the classical level.
+By making use of the equation of motion for the auxiliary field $B_{ij}$, we can fix the Lagrangian multiplier $\varLambda^{ij}$ as
 $\varLambda^{ij}=\frac{1}{N\sqrt{h}}\frac{\delta S}{\delta B_{ij}}.$ (5)
-Finally, by plugging the solution for $\varLambda^{ij}$ into the action (4), we can recast it in the form
+Finally, by plugging the solution of $\varLambda^{ij}$, the action (4) can be recast in the form
 $\tilde{S}=S\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}x\frac{\delta S}{\delta B_{ij}}\left(K_{ij}-B_{ij}\right).$ (6)
@@
 $\varPhi^{I}\coloneqq\left\{N^{i},\phi,B_{ij},N,h_{ij}\right\},$ (7)
-and the corresponding set of conjugate momenta is denoted by
+and the set of their conjugate momenta is denoted by
 $\varPi_{I}\coloneqq\left\{\pi_{i},p,p^{ij},\pi,\pi^{ij}\right\}.$ (8)
@@
 $\varPi_{I}$ $\coloneqq\frac{\delta\tilde{S}}{\delta\dot{\Phi}^{I}}.$ (9)
-Since all velocities $\dot{\varPhi}^{I}$ in the Lagrangian (6) appear linearly, the conjugate momenta automatically yield constraints among the phase space variables. Indeed, the conjugate momenta are
+Since all the velocities $\dot{\varPhi}^{I}$ in the Lagrangian (6) are linearized, the conjugate momenta autonatically correspond to constraints among phase space variables.
+Indeed, the conjugate momenta are
 $\left\{\pi_{i},p,p^{ij},\pi,\pi^{ij}\right\}$ $=$ $\left\{\frac{\delta\tilde{S}}{\delta\dot{N}^{i}},\frac{\delta\tilde{S}}{\delta\dot{\phi}},\frac{\delta\tilde{S}}{\delta\dot{B}_{ij}},\frac{\delta\tilde{S}}{\delta\dot{N}},\frac{\delta\tilde{S}}{\delta\dot{h}_{ij}}\right\}$ (10)
 $=$ $\left\{0^{i},0,0_{ij},0,\frac{1}{2N}\frac{\delta S}{\delta B_{ij}}\right\},$
-which lead to the primary constraints
+which result in the primary constraints
 $\pi_{i}\approx 0,\quad p\approx 0,\quad p^{ij}\approx 0,\quad\pi\approx 0,\quad\tilde{\pi}^{ij}\equiv\pi^{ij}-\frac{1}{2N}\frac{\delta S}{\delta B_{ij}}\approx 0.$ (11)
-Here and throughout this work, the symbol $\approx$” denotes “weak equality”, i.e., an equality that holds only when the constraints are satisfied.
-By contrast, “$=$” denotes strong equality, which holds regardless of whether constraints are imposed.
+Here and throughout this work $\approx$” represents “weak equality”, i.e., an equality that holds only when the constraints are satisfied.
+On the other hand, the strong equality “==” holds no matter the constraints are satisfied or not.
 We denote the set of primary constraints as
 $\varphi^{I}\coloneqq\left\{\pi_{i},p,p^{ij},\pi,\tilde{\pi}^{ij}\right\},$ (12)
-and define the subspace of the phase space where the primary constraints are satisfied as $\varGamma_{\mathrm{P}}$, which can be regarded as a hypersurface in the original phase space $\left\{\varPhi^{I},\varPi_{I}\right\}$.
-The canonical Hamiltonian is obtained via the Legendre transformation:
+and the subspace of the phase space when the primary constraints are satisfied as $\varGamma_{\mathrm{P}}$, which can be viewed as a hypersurface in the original phase space $\left\{\varPhi^{I},\varPi_{I}\right\}$.
+The canonical Hamiltonian is obtained by performing the Legendre transformation
 $H_{\mathrm{C}}$ $=\int\mathrm{d}^{3}x\left(\sum_{I}\varPi_{I}\dot{\varPhi}^{I}-N\sqrt{h}\tilde{\mathcal{L}}\right)\approx\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}},$ (13)
@@
 $X[\vec{\xi}]$ $\coloneqq$ $\int\mathrm{d}^{3}x\,\varPi_{I}\pounds_{\vec{\xi}}\varPhi^{I}$ (16)
-$=$ $\int\mathrm{d}^{3}x\left(\pi_{i}\pounds_{\vec{\xi}}N^{i}+p\pounds_{\vec{\xi}}\phi+p^{ij}\pounds_{\vec{\xi}}B_{ij}+\pi\pounds_{\vec{\xi}}N+\pi^{ij}\pounds_{\vec{\xi}}h_{ij}\right).$
-In the following, we briefly summarize the properties of $X[\vec{\xi}]$ as discussed in [68].
-Suppose $Q\left(\vec{x}\right)$ is a general spatial tensor field or density (with spatial indices suppressed for brevity) constructed from the phase space variables $\left\{\varPhi^{I},\varPi_{I}\right\}$, and let $\mathcal{F}$ be a scalar functional defined by
+$=$ $\int\mathrm{d}^{3}x\left(\pi_{i}\pounds_{\vec{\xi}}N^{i}+p\pounds_{\vec{\xi}}\phi+p^{ij}\pounds_{\vec{\xi}}B_{ij}+\pi\pounds_{\vec{\xi}}N+\pi^{ij}\pounds_{\vec{\xi}}h_{ij}\right),$
+In the following we briefly mention the properties of $X[\vec{\xi}]$ following [68].
+Supposing $Q\left(\vec{x}\right)$ is a general spatial tensor field or density (with spatial indices suppressed for briefness) built of phase space variables $\left\{\varPhi^{I},\varPi_{I}\right\}$, and $\mathcal{F}$ is a scalar functional defined by
 $\mathcal{F}\left[f,\varPhi^{I},\varPi_{I}\right]=\int\mathrm{d}^{3}x\,f(\vec{x})Q(\vec{x}),$ (17)
-where $f(\vec{x})$ is a smooth test function.
-Note that $f(\vec{x})$ may also carry appropriate spatial indices (suppressed here) such that $\mathcal{F}$ is invariant under time-independent spatial diffeomorphisms.
-The Poisson bracket between $X[\vec{\xi}]$ and $\mathcal{F}$ is given by
+with $f\left(\vec{x}\right)$ is a smoothing function.
+Note $f(\vec{x})$ also constains appropriate spatial indices (suppressed as well) such that $\mathcal{F}$ is invariant under time-independent spatial diffeomorphism.
+The Poisson bracket of $X[\vec{\xi}]$ and $\mathcal{F}$ are
 $[X[\vec{\xi}],\mathcal{F}]$ $=$ $X\left[[\vec{\xi},\mathcal{F}]\right]+\int\mathrm{d}^{3}x\frac{\delta\mathcal{F}}{\delta f}\pounds_{\vec{\xi}}f$ (18)
 $=$ $\int\mathrm{d}^{3}x\,Q(\vec{x})\pounds_{\vec{\xi}}f(\vec{x}).$
-In (18), the Poisson bracket $[\mathcal{F},\mathcal{G}]$ is defined as
-$[\mathcal{F},\mathcal{G}]\coloneqq\int\mathrm{d}^{3}z\left(\frac{\delta\mathcal{F}}{\delta\varPhi_{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPi^{I}(\vec{z})}-\frac{\delta\mathcal{F}}{\delta\varPi^{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPhi_{I}(\vec{z})}\right).$ (19)
-A key point is that the right-hand side of the second line in (18) vanishes on the constraint surface if $Q(\vec{x})$ is a constraint.
-By integrating by parts, $X[\vec{\xi}]$ can be written in a more convenient form:
+In (18), the Poisson bracket $\mathcal{F},\mathcal{G}]$ is defined by
+$\mathcal{F},\mathcal{G}]\coloneqq\int\mathrm{d}^{3}z\left(\frac{\delta\mathcal{F}}{\delta\varPhi_{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPi^{I}(\vec{z})}-\frac{\delta\mathcal{F}}{\delta\varPi^{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPhi_{I}(\vec{z})}\right).$ (19)
+The key point is that the right side of the second line of the equation vanishes on the constraint surface if $Q(\vec{x})$ is a constraint.
+Through integrations by parts, $X[\vec{\xi}]$ can be written in a more convenient form
 $X[\vec{\xi}]=\int\mathrm{d}^{3}x\,\xi^{i}\mathcal{C}_{i},$ (20)
@@
 $-2\sqrt{h}\mathrm{D}_{j}\left(\frac{p^{jk}}{\sqrt{h}}B_{ik}\right)+\pi_{j}\mathrm{D}_{i}N^{j}+\sqrt{h}\mathrm{D}_{j}\left(\frac{1}{\sqrt{h}}\pi_{i}N^{j}\right).$
-As we shall see later, the constraints $\mathcal{C}_{i}\approx 0$ are secondary constraints associated with the primary constraints $\pi_{i}\approx 0$.
-Furthermore, the Poisson bracket between $\mathcal{C}_{i}$ and any constraint $Q$ is given by
+As we shall see later, $\mathcal{C}_{i}\approx 0$ are actually the secondary constraints associated with the primary constraints $\pi_{i}\approx 0$.
+Moreover, the Poisson bracket of $\mathcal{C}_{i}$ with any constraint $Q$ can be obtained by
 $\left[X[\vec{\xi}],\mathcal{F}\right]=\int\mathrm{d}^{3}x\mathrm{d}^{3}y\,\xi^{i}(x)f(y)\left[\mathcal{C}_{i}(\vec{x}),Q(\vec{y})\right].$ (22)
@@
 $\left[\mathcal{C}_{i}(\vec{x}),Q(\vec{y})\right]\approx 0,\quad\text{for any}\quad Q\approx 0.$ (23)
-Due to the presence of primary constraints, the time evolution is governed by the total Hamiltonian, defined as
+Due to the presence of primary constraints, the time evolution is determined by the total Hamiltonian defined by
 $H_{\mathrm{T}}\coloneqq H_{\mathrm{C}}+\int\mathrm{d}^{3}y\sum_{I}\lambda_{I}(\vec{y})\varphi^{I}(\vec{y}),$ (24)
-where $\lambda_{I}:=\left\{v^{i},v,v_{ij},\lambda,\lambda_{ij}\right\}$ are undetermined Lagrange multipliers.
-Since the canonical Hamiltonian $H_{\mathrm{C}}$ is only well-defined on the subspace defined by the primary constraints $\varGamma_{\mathrm{P}}$, we may directly use $\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}$ given in (14) in place of $H_{\mathrm{C}}$ in (13) for subsequent calculations.
+where $\lambda_{I}:=\left\{v^{i},v,v_{ij},\lambda,\lambda_{ij}\right\}$ are the undetermined Lagrange multipliers.
+Since the canonical Hamiltonian $H_{\mathrm{C}}$ is well-defined only on the subspace defined by the primary constraints $\varGamma_{\mathrm{P}}$, we can directly use $\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}$ given in (14) instead of $H_{\mathrm{C}}$ in (13) in the subsequent calculations.
 III.2 Consistency condition
-Constraints must be preserved under time evolution. For the primary constraints, we require that
+Constraints must be preserved in time evolution. For the primary constraints, we must require that
 $\int\mathrm{d}^{3}y\left[\varphi^{I}(\vec{x}),\varphi^{J}(\vec{y})\right]\lambda_{J}(\vec{y})+\left[\varphi^{I}(\vec{x}),H_{\mathrm{C}}\right]\approx 0,$ (25)
-which are the so-called consistency conditions for the primary constraints. These conditions may yield further constraints.
-Without imposing additional assumptions on the structure of the Lagrangian in (6), the Poisson brackets between the primary constraints have been computed in detail in [85], and are summarized in Appendix B.
-For later convenience, we rewrite (25) in matrix form:
+which are the so-called consistency conditions for the primary constraints. These conditions may yields further constraints.
+With no further requirement on the structure of the Lagrangian in (6), the Poisson brackets between primary constraints have been calculated in detail in [85], which are summarized in Appendix B.
+For later convenience, we write (25) in matrix form
 $\int\mathrm{d}^{3}y\left(\begin{array}[]{ccccc}0&0&0&0&0\\ 0&0&0&0&\left[p(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&0&0&0&\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&0&0&0&\left[\pi(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&\left[\tilde{\pi}^{ij}(\vec{x}),p(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),p^{kl}(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),\pi(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\end{array}\right)\left(\begin{array}[]{c}v^{i}\\ v\\ v_{ij}\\ \lambda\\ \lambda_{ij}\end{array}\right)\approx\left(\begin{array}[]{c}-\mathcal{C}_{i}(\vec{x})\\ \frac{\delta S}{\delta\phi(\vec{x})}\\ 0\\ -\mathcal{C}(\vec{x})\\ \left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]\end{array}\right),$ (26)
 where $\mathcal{C}_{i}$ is defined in (22).
-Note that $\mathcal{C}_{i}$ must have vanishing Poisson brackets with all other constraints due to the property (22).
+Note $\mathcal{C}_{i}$ must have vanishing Poisson brackets with any other constraints due to the property (22).
 In the fourth line of (26), $\mathcal{C}$ is defined by
@@
 $\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]=\frac{\delta S}{\delta h_{ij}(\vec{x})}-\frac{1}{N(\vec{x})}\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{z})}N(\vec{z})B_{kl}(\vec{z}).$ (28)
-According to (129), $\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ in the third line of (26) is proportional to the kinetic matrix
+According to (129), $\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ in the third line in (26) is proportional to the kinetic matrix
 $\frac{\delta^{2}S_{B}}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})},$ (29)
 which we assume to be non-degenerate.
-This is because $B_{ij}$ is the auxiliary field corresponding to $K_{ij}$.
-The non-degeneracy of (29) implies that the kinetic term for $\dot{h}_{ij}$ is non-degenerate, which guarantees that the theory includes General Relativity as a limiting case, and in particular, allows the existence of gravitational waves.
-As a result, the Lagrange multipliers $\lambda_{ij}$ can be fixed from the third line of (26) as
+This is because $B_{ij}$ is the auxiliary field corresponding to $K_{ij}$, The non-degeneracy of (29) implies that the kinetic term for $\dot{h}_{ij}$ is not degenerate, which guarantees the theory includes General Relativity as its limit, and in particular, the existence of gravitational waves.
+As a result, the Lagrange multiplies $\lambda_{ij}$ can be fixed from the third line in (26) to be
 $\lambda_{ij}=0.$ (30)
-Then the second and fourth lines of (26) yield two additional secondary constraints:
+Then the second and the fourth line of (26) yield another two secondary constraints
 $\frac{\delta S}{\delta\phi(\vec{x})}\approx 0,\quad\mathcal{C}(\vec{x})\approx 0.$ (31)
-The constraints $\frac{\delta S}{\delta\phi(\vec{x})}$ and $\mathcal{C}(\vec{x})$ do not generate further constraints via consistency conditions, as will become clear in the next section.
-Having identified all the primary and secondary constraints, we are now in a position to classify them.
-According to the terminology of the Dirac-Bergmann algorithm, a constraint that has vanishing Poisson brackets with all other constraints (including itself) is called a first-class constraint; otherwise, it is second-class.
-The distinction is that each first-class constraint eliminates one degree of freedom, whereas a second-class constraint removes one phase space dimension, i.e., half a degree of freedom.
-From the discussion above on $\mathcal{C}_{i}$ and the constraint algebra in Appendix B, the Poisson brackets among all the constraints can be summarized in the so-called Dirac matrix:
+Constraints $\frac{\delta S}{\delta\phi(\vec{x})}$ and $\mathcal{C}(\vec{x})$ do not generate additional constraints due to consistency conditions, which will be clear in the next section.
+After finding all the primary and secondary constraints, we are now ready to classify all the constraints.
+According to the terminology of Dirac-Bergmann algorithm, a constraint that has vanishing Poisson brackets with all contraints (including itself) is dubbed a first-class constraint, otherwise a second-class constraint.
+The difference is that each first-class constraint eliminates a single degree of freedom, while a second-class constraint, which reduce one phase space dimension, eliminate half degree of freedom.
+According to the above discussion on $\mathcal{C}_{i}$ and constraint algebra in Appendix B, the Poisson brackets among all the constraints can be summarized in the so-called Dirac matrix:
 $\begin{array}[]{|c|cccccccc|}\hline\cr[\cdot,\cdot]&\pi_{i}(\vec{y})&\mathcal{C}_{i}(\vec{y})&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\tilde{\pi}^{kl}(\vec{y})&\mathcal{C}(\vec{y})&\frac{\delta S}{\delta\phi(\vec{y})}\\ \hline\cr\pi_{i}(\vec{x})&0&0&0&0&0&0&0&0\\ \mathcal{C}_{i}(\vec{x})&0&0&0&0&0&0&0&0\\ p(\vec{x})&0&0&0&0&0&X&X&X\\ \pi(\vec{x})&0&0&0&0&0&X&X&X\\ p^{ij}(\vec{x})&0&0&0&0&0&X&X&X\\ \tilde{\pi}^{ij}(\vec{x})&0&0&X&X&X&X&X&X\\ \mathcal{C}(\vec{x})&0&0&X&X&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&0&0&X&X&X&X&X&0\\ \hline\cr\end{array},$ (32)
-where an “$X$” denotes a generically non-vanishing Poisson bracket. In total, there are 22 constraints, which can be divided into two classes:
+where a “XX” stands for generally non-vanishing Poisson brackets. In total there are 22 constraints which can be divided into two classes:
 $\begin{array}[]{rl}6\text{ first-class: }&\pi_{i},\quad\mathcal{C}_{i},\\ 16\text{ second-class: }&p,\quad p^{ij},\quad\pi,\quad\tilde{\pi}^{ij},\quad\frac{\delta S}{\delta\phi},\quad\mathcal{C}.\end{array}$
-The number of degrees of freedom (DOF) is thus
-#$\#_{\mathrm{DOF}}$ $=$ $\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (33)
+The number of DOF is thus
+$\#_{\mathrm{DOF}}$ $=$ $\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (33)
 $=$ $\frac{1}{2}(2\times 17-2\times 6-16)$
-$=$ 3 3,3 ,
-which explicitly shows that a spatially covariant gravity theory where temporal derivatives arise only through the extrinsic curvature has 3 propagating degrees of freedom, regardless of the presence of additional auxiliary field(s).
+$=$ $3,$
+which explicit shows that spatially covariant gravity with temporal derivative arises only through the extrinsic curvature, has 3 propagating degrees of freedom, no matter with or without extra auxiliary field(s).
 IV Degeneracy Analysis
-To reduce the number of DOFs, additional conditions must be imposed on the Lagrangian (6).
-As we will see in this section, two conditions are required in order for the theory to propagate only two DOFs.
+To reduce the number of DOFs, additional conditions must be put on the Lagrangian (6). As we will see in this section, two conditions are needed in order to make the theory propagate only two DOFs.
 IV.1 The first condition
-As first-class constraints, $\pi_{i}$ and $\mathcal{C}_{i}$ correspond to spatial invariance.
-Without breaking spatial invariance, $\pi_{i}$ and $\mathcal{C}_{i}$ will always remain first-class constraints, which have vanishing Poisson brackets with all other constraints.
-For simplicity, we are allowed to omit these two first-class constraints in the subsequent degeneracy analysis.
-We therefore define a simplified Dirac matrix by omitting the two columns and rows corresponding to $\pi_{i}$ and $\mathcal{C}_{i}$:
+As being the first-class constraints, $\pi_{i}$ and $\mathcal{C}_{i}$ correspond to the spatial invariance.
+Without breaking the spatial invariance, $\pi_{i}$ and $\mathcal{C}_{i}$ will always be kept as the first-class constraints, which have vanishing Poisson brackets with any other constraints.
+For the sake of simplicity, we are allowed to omit these two first-class constraints in the subsequent degeneracy analysis.
+We thus define a simplified Dirac matrix by omitting the two columns and rows corresponding to $\pi_{i}$ and $\mathcal{C}_{i}$:
 $\mathcal{M}^{ab}(\vec{x},\vec{y})=\begin{array}[]{|c|ccc|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\tilde{\pi}^{kl}(\vec{y})&\mathcal{C}(\vec{y})&\frac{\delta S}{\delta\phi(\vec{y})}\\ \hline\cr p(\vec{x})&0&0&0&X&X&X\\ \pi(\vec{x})&0&0&0&X&X&X\\ p^{ij}(\vec{x})&0&0&0&X&X&X\\ \hline\cr\tilde{\pi}^{ij}(\vec{x})&X&X&X&X&X&X\\ \mathcal{C}(\vec{x})&X&X&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X&X&X&0\\ \hline\cr\end{array}.$ (34)
-Our goal is to find conditions that eliminate one scalar-type degree of freedom.
-Generally, we have two options: one is to introduce extra conditions to convert second-class constraints into first-class constraints, and the other is to generate new secondary constraints.
-In both cases, a necessary condition is that the simplified Dirac matrix $\mathcal{M}^{ab}$ must be degenerate.
+Our purpose is to find conditions to eliminate one scalar type degree of freedom.
+Generally we have two choices.
+One is to introduce extra conditions to change the second-class constraints into the first-class constraints, the other is to generate new secondary constraints.
+In both cases, a necessary conditions is that the simplified Dirac matrix $\mathcal{M}^{ab}$ must be degenerate.
 The determinant of the simplified Dirac matrix $\mathcal{M}^{ab}$ is
 $\det\mathcal{M}^{ab}(\vec{x},\vec{y})=\det\mathcal{B}^{ab}(\vec{x},\vec{y})\times\det\mathcal{B}^{ba}(\vec{y},\vec{x}),$ (35)
-where the matrix $\mathcal{B}^{ab}$ is the lower-left corner of $\mathcal{M}^{ab}$.
+where the matrix $\mathcal{B}^{ab}$ is the lower left corner of $\mathcal{M}^{ab}$.
 The degeneracy of $\mathcal{M}^{ab}$, i.e., $\det\mathcal{M}^{ab}=0$, thus implies that the sub-matrix $\mathcal{B}^{ab}$ of $\mathcal{M}^{ab}$ is degenerate, i.e., $\det\mathcal{B}^{ab}=0$.
-At first glance, the determinant of the matrix $\mathcal{B}^{ab}$ may appear to be the matrix form of the first 2DOF condition.
-However, this is not the case.
-The 2DOF condition should be an equation satisfied by the action and should not contain conjugate momenta.
-Otherwise, we obtain a relation between the action and the conjugate momenta, which is in fact a constraint.
-Upon careful inspection, all elements of the matrix $\mathcal{B}^{ab}$ are independent of the conjugate momenta, except for $\mathcal{B}^{21}$, given by
+Apparently, the determinant of matrix $\mathcal{B}^{ab}$ is the matrix form of the first 2DOF condition.
+This is not the case.
+Because the 2DOF condition should be an equation that the action satisfies, and should not contain conjugate momentum.
+Otherwise, we get the relationship between the action and the conjugate momentum, which is actually a constraint.
+By careful observation, all elements of the matrix $\mathcal{B}^{ab}$ are independent of conjugate momentum, except of $\mathcal{B}^{21}$ given by
 $\mathcal{B}^{21}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}(\vec{x}),p^{kl}(\vec{y})\right]$ (36)
 $=$ $2\delta^{3}(\vec{x}-\vec{y})\pi^{kl}(\vec{x})-\frac{\delta^{2}S}{\delta N(\vec{x})\delta B_{kl}(\vec{y})},$
-which can be rewritten as
+which can be transformed into
 $\mathcal{B}^{21}(\vec{x},\vec{y})\approx-\delta^{3}(\vec{x}-\vec{y})\frac{1}{N(\vec{y})}\frac{\delta S}{\delta B_{ij}(\vec{y})}+\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta N(\vec{y})}$ (37)
 on the constraint surface of $\tilde{\pi}^{ij}$.
-In subsequent calculations, we will always use (37) as the value of $\mathcal{B}^{21}$ instead of (36).
-As a result, the matrix form of the first 2DOF condition becomes
+In subsequent calculations, we always use (37) as the value of $\mathcal{B}^{21}$ instead of (36).
+As a result, the matrix form of the first 2DOF condition is
 $\mathcal{S}_{1}(\vec{x},\vec{y})=\det\mathcal{B}^{ab}(\vec{x},\vec{y})\approx 0,$ (38)
@@
 $\mathcal{B}^{ab}=\begin{array}[]{|c|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})\\ \hline\cr-2N(\vec{x})\tilde{\pi}^{ij}(\vec{x})&\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})}&N(\vec{x})\frac{\delta}{\delta N(\vec{y})}\left(\frac{1}{N(\vec{x})}\frac{\delta S}{\delta B_{ij}(\vec{x})}\right)&\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})}\\ -\mathcal{C}(\vec{x})&\frac{\delta^{2}S}{\delta N(\vec{x})\delta\phi(\vec{y})}&\frac{\delta^{2}S}{\delta N(\vec{x})\delta N(\vec{y})}&N(\vec{y})\frac{\delta}{\delta N(\vec{x})}\left(\frac{1}{N(\vec{y})}\frac{\delta S}{\delta B_{kl}(\vec{y})}\right)\\ \frac{\delta S}{\delta\phi(\vec{x})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta N(\vec{y})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta B_{kl}(\vec{y})}\\ \hline\cr\end{array}.$ (39)
-In the above, we have adjusted the coefficients of $\tilde{\pi}^{ij}$ and $\mathcal{C}$ in the matrix $\mathcal{B}^{ab}$ to make it more symmetric.
-Since $\mathcal{B}^{ab}$ is a degenerate matrix, there must exist a null eigenvector $\mathcal{V}_{a}$ satisfying
+In the above, we have adjusted the coefficients of $\tilde{\pi}^{ij}$ and $\mathcal{C}$ in the matrix $\mathcal{B}^{ab}$ to make the it more symmetric.
+Since $\mathcal{B}^{ab}$ is a degenerate matrix, there must be a null eigenvector $\mathcal{V}_{a}$ satisfying
 $\int\mathrm{d}^{3}y\,\mathcal{B}^{ab}(\vec{x},\vec{y})\mathcal{V}_{b}(\vec{y})\approx 0^{a}.$ (40)
@@
 $\mathcal{V}_{a}(\vec{x})=\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x})=\int\mathrm{d}^{3}y\,U_{a}(\vec{x},\vec{y})V(\vec{y}).$ (41)
-In (41), $V(\vec{x})$ is an arbitrary function of the spatial coordinates, and $U_{a}(\vec{x},\vec{y})$ depends on the phase space variables.
-One method to obtain the explicit form of the null eigenvector $\mathcal{V}_{a}$ is to calculate the adjoint matrix of $\mathcal{B}^{ab}$, as detailed in Appendix C.
-Assuming that $\mathcal{B}_{2}^{*}$ does not vanish, we can use a linear combination of the null eigenvector $\mathcal{V}_{a}$ and the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ to define a new constraint $\mathcal{C}^{\prime}$,
+In (41), $V(\vec{x})$ is an arbitrary function of spatial coordinates and $U_{a}(\vec{x},\vec{y})$ depends on the phase space variables.
+One method to get the explicit form of the null eigenvector $\mathcal{V}_{a}$ is to calculate the adjoint matrix of matrix $\mathcal{B}^{ab}$, which can be found in Appendix C.
+Assuming that $\mathcal{B}_{2}^{*}$ does not vanish, we can use the linear combination of the null eigenvector $\mathcal{V}_{a}$ and constraint $\tilde{\pi}^{ij}$, $\mathcal{C}$, $\frac{\delta S}{\delta\phi}$ to define a new constraint $\mathcal{C}^{\prime}$,
 $\int\mathrm{d}^{3}x\,\mathcal{C}^{\prime}(\vec{x})V(\vec{x})=\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}\tilde{\pi}^{ij}&\mathcal{C}&\frac{\delta S}{\delta\phi}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x})\approx 0.$ (42)
-In (42), $V(\vec{x})$ is the arbitrary function from (41).
-The advantage of introducing the new constraint $\mathcal{C}^{\prime}$ is that its Poisson brackets with the constraints $p$, $\pi$, and $p^{kl}$ all vanish:
+In (42), $V(\vec{x})$ is the arbitrary function in (41).
+The merit of introducing the new constraint $\mathcal{C}^{\prime}$ is that the Poisson brackets between $\mathcal{C}^{\prime}$ and constraints $p$, $\pi$, $p^{kl}$ are all vanishing
 $\left[\mathcal{C}^{\prime}(\vec{x}),p(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),\pi(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),p^{kl}(\vec{y})\right]\approx 0.$ (43)
-With this new set of primary constraints, the consistency condition for the new constraint $\mathcal{C}^{\prime}$ becomes
-$\frac{\mathrm{d}\mathcal{C}^{\prime}(\vec{x})}{\mathrm{d}t}$ $=\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{T}}\right]$ (44)
+With this new set of primary constraints, the consistency conditions of the new constraint $\mathcal{C}^{\prime}$ reduce to
+$\frac{\mathrm{d}\mathcal{C}^{\prime}(\vec{x})}{\mathrm{d}t}$ $=\left[\mathcal{C}^{\prime}(\vec{x})(\vec{x}),H_{\mathrm{T}}\right]$ (44)
 $=\int\mathrm{d}^{3}y\left[\mathcal{C}^{\prime}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\lambda_{kl}(\vec{y})+\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\approx 0$
-where $\lambda_{ij}$ has been fixed to zero in (26). If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$, (44) yields a new secondary constraint $\Phi(\vec{y})$:
+where $\lambda_{ij}$ has been fixed to be zero in (26). If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$, (44) yields a new secondary constraint $\Phi(\vec{y})$:
 $\Phi(\vec{y})=\left[\mathcal{C}^{\prime}(\vec{y}),H_{C}\right]\approx 0.$ (45)
-Otherwise, if no secondary constraint arises, we can still find a new first-class constraint by taking a linear combination of the existing constraints $p$, $\pi$, and $p^{ij}$.
-The necessary and sufficient condition for generating secondary constraints is that the consistency condition is not automatically satisfied (instead of requiring the simplified Dirac matrix to be degenerate).
-Therefore, after imposing the consistency condition, two scenarios may arise:
+Otherwise, there is no secondary constraint, but we can always find a new first class constraint through linear combination of constraints $p$, $\pi$ and $p^{ij}$.
+Because the necessary and sufficient condition for generating secondary constraints is that the consistency condition cannot be automatically satisfied, rather than requiring the simplified Dirac matrix to be degenerate.
+Therefore, after requiring the consistency condition, two cases may arise:
   1. 1.
-    The consistency condition is not automatically satisfied, and an additional secondary constraint is needed to enforce it;
+    The consistency condition is not automatically satisfied, so that an additional secondary constraint is required to ensure that the consistency condition is satisfied;
   2. 2.
-    The consistency condition is automatically satisfied, and a first-class constraint can be constructed from a linear combination of existing constraints.
-When we impose a specific 2DOF condition, we must check whether it violates the consistency condition and leads to additional secondary constraints.
-Such a newly generated constraint must be taken into account in the subsequent analysis.
+    The consistency condition is automatically satisfied, and a first class constraint is obtained by combining the existing constraints.
+When we require a specific 2DOF condition, we should check whether this condition violates the consistency condition and generates additional secondary constraints.
+We should consider this newly generated secondary constraint in the subsequent discussion.
 IV.2 The second condition
@@
 $\int\mathrm{d}^{3}x\,\mathcal{\pi}^{\prime}(\vec{x})V(\vec{x})\equiv\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}p^{ij}&\pi&p\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}^{\prime}_{1}\right)_{ij}&\mathcal{\mathcal{V}}^{\prime}_{2}&\mathcal{\mathcal{V}}^{\prime}_{3}\end{array}\right)(\vec{x})\approx 0,$ (46)
-where again $V(\vec{x})$ is the arbitrary function in (41). Similar to the constraint $\mathcal{C}^{\prime}$, the new constraint $\mathcal{\pi}^{\prime}$ has vanishing Poisson brackets:
+where again $V(\vec{x})$ is the arbitrary function in (41). Similar to constraint $\mathcal{C}^{\prime}$, the new constraint $\mathcal{\pi}^{\prime}$ has the vanishing Poisson brackets
 $\left[\tilde{\pi}^{ij}(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]\approx\left[\frac{\delta S}{\delta\phi(\vec{x})},\mathcal{\pi}^{\prime}(\vec{y})\right]\approx 0.$ (47)
 By making use of the constraints $\mathcal{\pi}^{\prime}$, $\mathcal{C}^{\prime}$ and $\Phi$, the simplified Dirac matrix becomes
-=$\mathcal{M}^{\prime}{}^{ab}=\begin{array}[]{|c|ccccc|}\hline\cr[\cdot,\cdot]&\mathcal{\pi}^{\prime}(\vec{y})&\left\{p(\vec{y}),p^{kl}(\vec{y})\right\}&\mathcal{C}^{\prime}(\vec{y})&\left\{\tilde{\pi}^{kl}(\vec{y}),\frac{\delta S}{\delta\phi(\vec{y})}\right\}&\Phi(\vec{x})\\ \hline\cr\mathcal{\pi}^{\prime}(\vec{x})&0&0&0&0&X\\ \left\{p(\vec{x}),p^{kl}(\vec{x})\right\}&0&0&0&-\mathcal{C}^{ba}(\vec{y},\vec{x})&X\\ \mathcal{C}^{\prime}(\vec{x})&0&0&X&X&X\\ \left\{\tilde{\pi}^{kl}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{x})}\right\}&0&\mathcal{C}^{ab}(\vec{x},\vec{y})&X&X&X\\ \Phi(\vec{x})&X&X&X&X&X\\ \hline\cr\end{array}$ (48)
-Since the first 2DOF condition $\mathcal{S}_{1}(\vec{x},\vec{y})$ has been imposed, which yields a secondary constraint that reduces the dimension of the phase space by one and thus removes half a degree of freedom, we need to impose an additional 2DOF condition to reduce the number of degrees of freedom to two.
-This means that the simplified Dirac matrix must be degenerate again.
-The determinant of this matrix is proportional to $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]$, $\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]$, and $\det\mathcal{C}^{ab}(\vec{x},\vec{y})$.
-If the determinant of this matrix is required to vanish, two types of 2DOF conditions can be obtained:
+[…]
+Since the first 2DOF condition $\mathcal{S}_{1}(\vec{x},\vec{y})$ has been imposed, which yields a secondary constraint to reduced the dimension of the phase space by one and thus reduce a half degree of freedom.
+We need to require an additional 2DOF condition to reduce the number of degrees of freedom to 2, which means that the simplified Dirac matrix is required to be degenerate again.
+The determinant of the matrix is proportional to $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]$, $\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]$ and $\det\mathcal{C}^{ab}(\vec{x},\vec{y})$.
+If the determinant of this matrix is required to be vanishing, two types of 2DOF conditions can be obtained:
 $\mathcal{S}_{2}^{(1)}(\vec{x},\vec{y})$ $=$ $\det\mathcal{C}^{ab}(\vec{x},\vec{y})\times\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]=\det\mathcal{D}^{ab}(\vec{x},\vec{y})\approx 0,$ (49)
@@
 $\mathcal{D}^{ab}(\vec{x},\vec{y})$ $=\begin{array}[]{|c|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})\\ \hline\cr\tilde{\pi}^{ij}(\vec{x})&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X\\ \Phi(\vec{x})&X&X&X\\ \hline\cr\end{array}.$ (55)
-The 2DOF condition $\mathcal{S}_{2}^{(1)}$ will generate a secondary constraint if the consistency condition is not automatically satisfied.
+The 2DOF condition $\mathcal{S}_{2}^{(1)}$ will generate a secondary constraint, if the consistency condition is not automatically satisfied.
 On the other hand, the 2DOF condition $\mathcal{S}_{2}^{(2)}$ clearly generates a first-class constraint.
-If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]=0$ in the previous section, the second 2DOF conditions become
+If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]=0$ in the previous section, the second 2DOF conditions will become
 $\mathcal{S}^{\prime}{}_{2}^{(1)}(\vec{x},\vec{y})$ $=$ $\det\mathcal{C}^{ab}(\vec{x},\vec{y})\approx 0,$ (56)
 $\mathcal{S}^{\prime}{}_{2}^{(2)}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]\approx 0,$ (57)
-which are covered by the TTDOF conditions $\mathcal{S}_{2}^{(1)}$ and $\mathcal{S}_{2}^{(2)}$.
-The space of theories satisfying the 2DOF conditions $\mathcal{S}_{1}$, $\mathcal{S}{}_{2}^{(1)}$, and $\mathcal{S}_{2}^{(2)}$ can be divided into three branches:
+which is covered by TTDOF conditions $\mathcal{S}_{2}^{(1)}$and $\mathcal{S}_{2}^{(2)}$.
+The space of theories satisfying the 2DOF conditions $\mathcal{S}_{1}$, $\mathcal{S}{}_{2}^{(1)}$ and $\mathcal{S}_{2}^{(2)}$ can be divided into three branches:
   1. 1.
-    Linear dependence branch: If the secondary constraints are linearly dependent, the 2DOF conditions will be automatically satisfied.
-However, in this case, the number of degrees of freedom will not change.
-Due to the linear dependence, the number of secondary constraints is reduced by one, which offsets the half-degree-of-freedom reduction caused by the 2DOF conditions.
-For example, if the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ are linearly dependent, the 2DOF condition $\mathcal{S}_{1}$ is automatically satisfied.
-However, the number of degrees of freedom remains unchanged.
+    Linear constraints branch: If the secondary constraints are linearly dependent, 2DOF conditions will be strongly vanishing.
+However, in this case the number of degrees of freedom will not change.
+Due to the linear dependence, the number of secondary constraints will be reduced by one, which will offset the half degree of freedom reduction caused by the 2DOF conditions.
+For example, if constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$ are linearly dependent, 2DOF condition $\mathcal{S}_{1}$ is automatically satisfied.
+However, in this case the number of degrees of freedom remains unchanged.
   2. 2.
-    Non-physical branch: Some 2DOF conditions may render the theory trivial.
-For example, if the secondary constraints generated by the consistency condition still fail to satisfy the consistency condition, additional secondary constraints will be generated.
-These additional constraints may reduce the number of degrees of freedom to less than two, which is unphysical.
+    Non-physical branch: Some 2DOF conditions may make the theory be trival.
+For example, if the secondary constraints generated by the consistency condition still cannot satisfy the consistency condition, additional secondary constraints will be generated.
+However, these additional secondary constraints may reduce the number of degrees of freedom to be less than 2, which is an unphysical situation.
   3. 3.
-    Physical branch: In this branch, we obtain a scalar-tensor theory with only two tensorial degrees of freedom.
-In the following, when we refer to 2DOF conditions, we always mean those corresponding to the third physical branch.
+    Physical branch: In this branch, we get a scalar-tensor theory with only two tensorial degrees of freedom.
+In the following when we mention 2DOF conditions, we always refer to the third physical branch.
 V A concrete model of $d=2$
 V.1 The Lagrangian
-In [57], [58] (see also [74]), monomials of SCG have been constructed and classified according to the total number of derivatives $d$ in each monomial.
-As an example of our formalism, in this section we apply the two conditions to the concrete model with $d=2$.
-The action is given by
+In [57, 58] (see also [74]), monomials of SCG have been constructed and classified according to the total number of derivatives $d$ in each monomial.
+As an example of our formalism, in this section we will apply the two conditions to the concrete model of $d=2$.
+The action of is given by
 $\tilde{S}_{2}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right)=S_{2}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}x\frac{\delta S_{2}}{\delta B_{ij}}\left(K_{ij}-B_{ij}\right).$ (58)
-As discussed earlier, we introduce the variable $B_{ij}$ to make the explicit inversion of the velocity $\dot{h}_{ij}$ in terms of the momentum possible.
+As being discussed before, we introduce the variable $B_{ij}$ in order to make the explicit reversion of velocity $\dot{h}_{ij}$ in terms of the momentum possible.
 The equivalent action $S_{2}$ can be written as
@@
 $a_{i}=\frac{1}{N}\mathrm{D}_{i}N.$ (61)
-For later convenience, we evaluate some of the elements of $\mathcal{B}^{ab}$ as
+For later convenience, we can evaluate some of the elements of $\mathcal{B}^{ab}$ as
 $\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{mn}(\vec{y})}$ $=$ $2N\sqrt{h}\left(\frac{1}{2}c_{1}(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm})+c_{2}\delta^{ij}\delta^{mn}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}),$ (62)
@@
 $\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})}$ $=$ $2N\sqrt{h}\left(\frac{\delta c_{1}}{\delta\phi}B^{ij}+\frac{\delta c_{2}}{\delta\phi}B\delta^{ij}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}).$ (64)
-To prevent the dynamical terms in the theory from degenerating, we assume
+In order to prevent the dynamical terms of the theory from degenerating, we assume
 $\det\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{mn}(\vec{y})}=\left(2N\sqrt{h}\delta^{3}(\vec{x}-\vec{y})\right)^{6}c_{1}^{6}(c_{1}+3c_{2})\neq 0,$ (65)
@@
 As long as (66) is satisfied, the non-physical branch of the 2DOF conditions is naturally removed.
-Other elements of the matrix $\mathcal{B}^{ab}$ are evaluated as
+Other elements of matrix $\mathcal{B}^{ab}$ are evaluated to be
 $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}+\right$. (67)
@@
 $\quad\left.-\frac{d_{2}}{N}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
-where $\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}$ represents the terms in $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ that are independent of derivatives of the delta function.
-We use this notation because their explicit expressions are very complicated, and only the coefficients of the derivatives of the delta function are relevant to the subsequent analysis.
+where $\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}$ represents terms in $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ that are independent of derivatives of the delta function.
+We use this notation because their specific expressions are very complicated and only the coefficients of derivatives of the delta function are relevant to the subsequent analysis.
 V.2 The first condition
 We have already obtained the first 2DOF condition in (38).
-To get the specific form of the 2DOF conditions for this model, a straightforward approach is to calculate the determinant of the matrix $\mathcal{B}^{ab}$, which is, however, very complicated and not illuminating.
+To get the specific form of the 2DOF conditions for the this model, a straightforward approach is to calculate the determinant of the matrix $\mathcal{B}^{ab}$, which is however is very complicated and not illuminating.
 Alternatively, we can choose an equivalent approach by assuming that a null eigenvector $\mathcal{V}_{a}(\vec{x})$ exists and find the expression of this null eigenvector.
-In this approach, we can determine which conditions are required to make the Dirac matrix $\mathcal{B}^{ab}$ degenerate.
+In this approach, we can find which conditions are required to make the Dirac matrix $\mathcal{B}^{ab}$ degenerate.
 We assume $U_{a}(\vec{x},\vec{y})$ in (41) is independent of the derivative of the delta function.
-As a result, different orders of derivatives of the delta function will be separated in the matrix $\mathcal{B}^{ab}$,
+As a result, different orders of derivative of the delta function will be separated in the matrix $\mathcal{B}^{ab}$,
 $\mathcal{B}^{ab}(\vec{x},\vec{y})=\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})+\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})+\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y}),$ (71)
-in which terms with different orders (of derivatives of the delta function) in the matrix $\mathcal{B}^{ab}$ separately satisfy
+in which terms with different order (of derivative of the delta function) in the matrix $\mathcal{B}^{ab}$ separately satisfies
 $\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(i)}^{ab}(\vec{x},\vec{y})\approx 0^{b},\quad i=0,1,2.$ (72)
-Note that by observing the form of the secondary constraint $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$, (72) can only be weakly satisfied when $i=0$.
-In the matrices $\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})$ and $\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y})$, only the $\times 2$ 2 submatrices in the lower left corner are nonvanishing, which are
+Note by observing the form of the secondary constraint $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$, (72) can only be weakly satisfied when $i=0$.
+In matrix $\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})$ and $\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y})$, only the $2\times 2$ sub-matrices in the lower left corner are not vanishing, which are
 $\left(\begin{array}[]{cc}2\left(\frac{d_{1}}{N}+\frac{\partial d_{1}}{\partial N}-\frac{1}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathrm{D}^{i}\phi-\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}a^{i}&-\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathrm{D}^{i}\phi+2\left(\frac{c_{4}}{N^{2}}-\frac{1}{N}\frac{\partial c_{4}}{\partial N}\right)a^{i}\\ -2\frac{\partial d_{1}}{\partial\phi}\mathrm{D}^{i}\phi-2\left(N\frac{\partial d_{1}}{\partial N}+d_{1}\right)a^{i}&-2\left(\frac{\partial d_{1}}{\partial N}+\frac{d_{1}}{N}\right)\mathrm{D}^{i}\phi+\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)a^{i}\end{array}\right)_{x}\mathrm{D}_{x^{i}}\delta^{3}\left(\vec{x}-\vec{y}\right)$ (73)
@@
 respectively.
-In the above, a subscript “$x$” denotes functional dependence on $\vec{x}$.
+In the above a subscript “xx” denotes functional dependence on $\vec{x}$.
 In the subsequent discussion, we omit $\vec{x}$ for brevity when the dependence on $\vec{x}$ is clear from the context.
-According to (74), for the null eigenvalue equation of $\mathcal{B}_{(2)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ proportional to $\mathrm{D}^{x^{i}}\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
-$d_{2}\mathcal{V}_{2}+2Nd_{1}\mathcal{V}_{3}$ $=$ 0 0,0 , (75)
-$2c_{4}\mathcal{V}_{2}+Nd_{2}\mathcal{V}_{3}$ $=$ 0 0.0 . (76)
-As for the null eigenvalue equation of $\mathcal{B}_{(1)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ only related to $\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
-$\left(\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathcal{V}_{2}-2\frac{\partial d_{1}}{\partial\phi}\mathcal{V}_{3}$ $=$ 0 0,0 , (77)
-$\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2N\frac{\partial d_{1}}{\partial N}+2d_{1}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (78)
-$\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2\frac{\partial d_{1}}{\partial N}+\frac{2d_{1}}{N}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (79)
-$\left(\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\right)\mathcal{V}_{2}-\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)\mathcal{V}_{3}$ $=$ 0 0.0 .
-(80)
+According to (74), for the null eigenvalue equation of $\mathcal{B}_{(2)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ is proportional to $\mathrm{D}^{x^{i}}\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$, which yields
+$d_{2}\mathcal{V}_{2}+2Nd_{1}\mathcal{V}_{3}$ $=$ $0,$ (75)
+$2c_{4}\mathcal{V}_{2}+Nd_{2}\mathcal{V}_{3}$ $=$ $0.$ (76)
+As for the null eigenvalue equation of $\mathcal{B}_{(1)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ only related to $\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$, which yields
+$\left(\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathcal{V}_{2}-2\frac{\partial d_{1}}{\partial\phi}\mathcal{V}_{3}$ $=$ $0,$ (77)
+$\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2N\frac{\partial d_{1}}{\partial N}+2d_{1}\right)\mathcal{V}_{3}$ $=$ $0,$ (78)
+$\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2\frac{\partial d_{1}}{\partial N}+\frac{2d_{1}}{N}\right)\mathcal{V}_{3}$ $=$ $0,$ (79)
+$\left(\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\right)\mathcal{V}_{2}-\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)\mathcal{V}_{3}$ $=$ $0.$ (80)
 There are four equations since the coefficients of $\mathrm{D}^{i}\phi$ and $a^{i}$ need to be proportional to the null eigenvector.
-We assume that $d_{1}$, $d_{2}$, and $c_{4}$ are nonvanishing.
+We assume that $d_{1}$, $d_{2}$ and $c_{4}$ are not vanishing.
 We can simply choose
@@
 $\mathcal{V}_{3}$ $=$ $-d_{2},$ (82)
-and obtain two independent equations
-$d_{2}^{2}-4d_{1}c_{4}$ $=$ 0 0,0 , (83)
-$\left(\frac{d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)2Nd_{1}+\frac{\partial d_{1}}{\partial\phi}d_{2}$ $=$ 0 0.0 . (84)
-The other two equations are not independent and can be obtained by combining these two.
+and we can get two independent equations
+$d_{2}^{2}-4d_{1}c_{4}$ $=$ $0,$ (83)
+$\left(\frac{d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)2Nd_{1}+\frac{\partial d_{1}}{\partial\phi}d_{2}$ $=$ $0.$ (84)
+The other two equations are not independent and can be obtained by combining these two equations.
 We have a set of special solutions
@@
 where $A$ and $B$ are constants. As we will see, this set of special solutions will simplify the calculations in the subsequent discussion.
-When we discuss the degeneracy of the matrix $\mathcal{B}_{(0)}^{ab}$, we need to be careful to identify which parts are weakly equal to zero.
-To clarify this, we first discuss how secondary constraints participate in the weak equality.
-Since in the Lagrangian 58 the velocity term $\dot{h}_{ij}$ (encoded in $K_{ij}$) has been linearized, each canonical momentum corresponds to a primary constraint.
-Then all the secondary constraints can be combined into forms that are independent of canonical momenta.
+When we discuss the degeneracy of the matrix $\mathcal{B}_{(0)}^{ab}$, we need to be careful to identify which parts weakly equal to zero.
+To make this clear, we should first discuss how secondary constraints participate in the weakly equal equation.
+Since in the Lagrangian 58 the the velocity term $\dot{h}_{ij}$ (encoded in $K_{ij}$) has been linearized, each canonical momentum will correspond to a primary constraint.
+Then all the secondary constraints can be combined into a form that are independent of canonical momenta.
 For example, we make a linear combination of the constraints $\tilde{\pi}^{ij}$ and $\mathcal{C}$,
@@
 $\left.+\left(B\frac{\partial c_{2}}{\partial\phi}-A\Big(-\frac{c_{2}}{N}+\frac{\partial c_{2}}{\partial N}\Big)\right)B^{2}+\left(B\frac{\partial c_{3}}{\partial\phi}-A\Big(\frac{c_{3}}{N}+\frac{\partial c_{3}}{\partial N}\Big)\right)R\right],$
-which does not depend on the special solution (85) chosen before.
-Since we have replaced $\mathcal{B}^{21}$ with a term independent of the canonical momentum, (38) can be equivalently written as
+which does not depend on the special solution (85) we chose before.
+Since we have replaced $\mathcal{B}^{21}$ with a term which is independent of the canonical momentum, (38) can be equivalently written as
 $\det\mathcal{B}^{ab}(\vec{x},\vec{y})=\alpha(\vec{x},\vec{y})\mathcal{C}_{2}(\vec{x})+\beta(\vec{x},\vec{y})\frac{\delta S}{\delta\phi(\vec{x})}.$ (88)
@@
 we obtain
-$\left(\mathcal{V}_{1}\right)_{ij}=-2d_{1}N\mathcal{V}_{ij}^{(N)}+d_{2}\mathcal{V}_{ij}^{(\phi)},$ (90)
+$\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}=-2d_{1}N\mathcal{V}_{ij}^{(N)}+d_{2}\mathcal{V}_{ij}^{(\phi)},$ (90)
 where
@@
 $\mathcal{V}_{ij}^{(\phi)}$ $=$ $\frac{1}{c_{1}}\frac{\partial c_{1}}{\partial\phi}B_{ij}+\left(\frac{1}{c_{1}+3c_{2}}\frac{\partial c_{2}}{\partial\phi}-\frac{c_{2}}{(c_{1}+3c_{2})c_{1}}\frac{\partial c_{1}}{\partial\phi}\right)B\delta_{ij}.$ (92)
-Since we have previously chosen a set of special solutions for $d_{1}$, $d_{2}$, and $c_{4}$, equation (72) with $i=0$ can be written as
+Since we have previously chosen a set of special solutions for $d_{1}$, $d_{2}$ and $c_{4}$, equation (72) with $i=0$ can be written as
 $\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})=D_{1}^{b}(\vec{y})B_{ij}(\vec{y})B^{ij}(\vec{y})+D_{2}^{b}(\vec{y})B^{2}(\vec{y})+D_{3}^{b}(\vec{y})R(\vec{y}),\qquad b=1,2.$ (93)
-Since there are no terms such as $a_{i}a^{i}$ in equation (93), the coefficient $\beta(\vec{x},\vec{y})$ is fixed to zero. The above formula can equivalently be written as
+Since there are no terms such as $a_{i}a^{i}$ in equation (93), the coefficient $\beta(\vec{x},\vec{y})$ is fixed to zero. The above formula can be equivalently written as
 $\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})=\alpha^{b}(\vec{y})\mathcal{C}_{2}(\vec{y}),\qquad b=1,2.$ (94)
-In order for (94) to have a solution, the coefficients of $B_{ij}B^{ij}$, $B^{2}$, and $R$ must vanish. There will be 6 equations for $c_{1}$, $c_{2}$, and $c_{3}$. We assume that
+In order to make (94) have a solution, the coefficients of $B_{ij}B^{ij}$, $B^{2}$ and $R$ need to be vanishing. There will be 6 equations for $c_{1}$, $c_{2}$ and $c_{3}$. We assume that
 $\frac{c_{1}}{c_{2}}=\mathrm{const}.$ (95)
-This assumption simplifies the expression of $D_{2}^{b}$.
+This assumption will simplify the expression of $D_{2}^{b}$.
 By fixing the coefficients of $B_{ij}B^{ij}$ and $R$ in (94), we obtain 4 differential equations
-$A\left[\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{1}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 , (96)
-$A\left[\frac{2}{N}\frac{\partial c_{1}}{\partial N}+\frac{\partial^{2}c_{1}}{\partial N^{2}}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial N}-\alpha^{2}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{2}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 ,
-$A\left[\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{1}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0,0 ,
-$A\left[\frac{2}{N}\frac{\partial c_{3}}{\partial N}+\frac{\partial^{2}c_{3}}{\partial N^{2}}-\alpha^{2}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{2}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0.0 .
-(97)
-As long as equations (95) and (96)-(97) are satisfied, the coefficient of $B^{2}$ will be automatically zero.
+$A\left[\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{1}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big(\frac{\partial c_{1}}{\partial\phi}\Big)^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ $0,$ (96)
+$A\left[\frac{2}{N}\frac{\partial c_{1}}{\partial N}+\frac{\partial^{2}c_{1}}{\partial N^{2}}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial N}-\alpha^{2}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{2}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ $0,$
+$A\left[\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{1}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ $0,$
+$A\left[\frac{2}{N}\frac{\partial c_{3}}{\partial N}+\frac{\partial^{2}c_{3}}{\partial N^{2}}-\alpha^{2}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{2}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ $0.$ (97)
+As long as equations (95) and (96)-(97) are satisfied, the coefficient of $B^{2}$ will be automatically fixed to be zero.
 Therefore, we do not have to discuss the differential equation corresponding to the coefficient of $B^{2}$ here.
@@
 $c_{3}$ $=$ $\frac{D}{N}g(\phi),$ (100)
-where the coefficients $C$ and $D$ are both constants.
-By plugging the expressions of $c_{1}$ and $c_{3}$ into (96)-(97), we can simplify the four differential equations as
-$\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (101)
-$\alpha^{2}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (102)
-$\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0,0 , (103)
-$\alpha^{2}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0.0 . (104)
+where the coefficients $C$ and $D$ are both constant.
+By plugging the expressions of $c_{1}$ and $c_{3}$ in (96)-(97), we can simplify the four differential equations
+$\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big(\frac{\partial c_{1}}{\partial\phi}\Big)^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}$ $=$ $0,$ (101)
+$\alpha^{2}\frac{\partial c_{1}}{\partial\phi}$ $=$ $0,$ (102)
+$\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}$ $=$ $0,$ (103)
+$\alpha^{2}\frac{\partial c_{3}}{\partial\phi}$ $=$ $0.$ (104)
 Finally, constraint $\mathcal{C}_{2}$ becomes
 $\mathcal{C}_{2}$ $=BN\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B_{ij}B^{ij}+\frac{\partial c_{2}}{\partial\phi}B^{2}+\frac{\partial c_{3}}{\partial\phi}R\right).$ (105)
-As we can see, if both $c_{1}$ and $c_{3}$ are independent of $\phi$, constraint $\mathcal{C}_{2}$ will not exist.
-The number of second-class constraints will be reduced by one, and the number of degrees of freedom will increase by 1/2.
-This is the linear-constraints branch of the 2DOF condition mentioned in the previous section, which is not the case we are interested in.
-Therefore, we assume that $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously.
-As a result, the coefficient $\alpha^{2}$ must be fixed to zero.
+As we can see, if both $c_{1}$ and $c_{3}$ are independent of $\phi$, constraint $\mathcal{C}_{2}$ will do not exist.
+The number of second-class constraints will be reduced by one, and the number of degrees of freedom will be increased by 1/2.
+This is the linear constraints branch of 2DOF condition we mentioned in the previous section, which is not the case we are interested in.
+Therefore we assume that $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously.
+As a result, the coefficient $\alpha^{2}$ has to be fixed to be zero.
 V.3 The second condition
-According to the above discussion, after imposing the first 2DOF condition $\mathcal{S}_{1}$, we can obtain the constraint $\mathcal{C}^{\prime}$ from the null eigenvector $\mathcal{V}_{a}(\vec{x})$,
+According to the above discussion, after requiring the first 2DOF condition $\mathcal{S}_{1}$, we can obtain constraint $\mathcal{C}^{\prime}$ by the null eigenvector $\mathcal{V}_{a}(\vec{x})$,
 $\int\mathrm{d}^{3}x\,\mathcal{C}^{\prime}(\vec{x})V(\vec{x})\equiv\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}-2N\tilde{\pi}^{ij}&-\mathcal{C}&\frac{\delta S}{\delta\phi}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x}),$ (106)
@@
 $\mathcal{C}^{\prime}=BN\left[\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B^{ij}B_{ij}+\Big(\frac{2c_{2}}{c_{1}}\frac{\partial c_{1}}{\partial\phi}-\frac{\partial c_{2}}{\partial\phi}\Big)B^{2}-\frac{\partial c_{3}}{\partial\phi}R\right)-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial\phi}\pi^{ij}B_{ij}\right].$ (107)
-The Poisson brackets between the constraint $\mathcal{C}^{\prime}$ and the constraints $p$, $\pi$, and $p^{kl}$ are
+Poisson brackets between the constraint $\mathcal{C}^{\prime}$ and constraints $p$, $\pi$, $p^{kl}$ are
 $\left[\mathcal{C}^{\prime}(\vec{x}),p(\vec{y})\right]$ $=\left[-BN\frac{\partial}{\partial\phi}\left(\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial\phi}\right)\tilde{\pi}^{ij}B_{ij}-\frac{\partial c_{3}}{\partial\phi\partial\phi}\left(\frac{\partial c_{3}}{\partial\phi}\right)^{-1}\mathcal{C}_{2}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (108)
@@
 $\left[\mathcal{C}^{\prime}(\vec{x}),p^{kl}(\vec{y})\right]$ $=-\frac{BN(\vec{x})}{c_{1}(\vec{x})}\frac{\partial c_{1}(\vec{x})}{\partial\phi(\vec{x})}\tilde{\pi}^{kl}(\vec{x})\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (110)
-which all vanish on the constraint surface. The Poisson bracket between the constraint $\mathcal{C}^{\prime}$ and itself is generally nonzero, unless
+which all vanish on the constraint surface. Poisson bracket between constraint $\mathcal{C}^{\prime}$ and itself is generally nonzero, unless
 $\frac{\partial c_{1}}{\partial\phi}=0,\quad\text{or}\quad\frac{\partial c_{3}}{\partial\phi}=0.$ (111)
-As mentioned above, $c_{1}$ and $c_{3}$ cannot both be independent of $\phi$ simultaneously; otherwise, the number of degrees of freedom will increase by 1/21/21 / 2.
-If we require the second 2DOF condition $\mathcal{S}_{2}^{(2)}$ to hold, there are two cases.
-In the first case, if we choose $\frac{\partial c_{1}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
+As we have mentioned above, $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously, otherwise the number of degrees of freedom will increased by $1/2$.
+If we require the second 2DOF condition $\mathcal{S}_{2}^{(2)}$ to be valid, there will be two cases.
+In the first case, if we choose $\frac{\partial c_{1}}{\partial\phi}=0$, constraint $\mathcal{C}_{2}$ will become
 $\mathcal{C}_{2}=BN\sqrt{h}\frac{\partial c_{3}}{\partial\phi}R\approx 0.$ (112)
-This case is clearly non-physical, as it fixes $R$ to vanish.
-On the other hand, if we choose $\frac{\partial c_{3}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
+This case is clearly non-physical as it fixes $R$ to be vanishing.
+On the other hand, if we choose $\frac{\partial c_{3}}{\partial\phi}=0$, constraint $\mathcal{C}_{2}$ will become
 $\mathcal{C}_{2}$ $=BN\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B_{ij}B^{ij}+\frac{\partial c_{2}}{\partial\phi}B^{2}\right)\approx 0.$ (113)
-In this case, we can solve for all the coefficients:
+In this case, we can solve all the coefficients
 $c_{1}$ $=Nf(\phi),\quad c_{2}=CNf(\phi),\quad c_{3}=\frac{D}{N},$
@@
 The resulting theory thus corresponds to a SCG with an auxiliary scalar field of $d=2$ that propagates two degrees of freedom.
-The constraint coefficients $\alpha^{i}$ in the previous differential equation are also uniquely determined:
+The constraint coefficients $\alpha^{i}$ in the previous differential equation are also uniquely determined
 $\alpha^{1}=-\left(\frac{\partial f(\phi)}{\partial\phi}\right)^{-1}\left(\frac{2}{f(\phi)}\Big(\frac{\partial f(\phi)}{\partial\phi}\Big)^{2}-\frac{\partial^{2}f(\phi)}{\partial\phi^{2}}\right),\quad\alpha^{2}=0.$ (115)
@@
 whose determinant is obviously zero.
-Based on the discussion of null eigenvectors in Appendix C, we can find a null eigenvector $\mathcal{V}_{a}^{(\mathcal{E})}$ to construct a new first-class constraint $\tilde{\mathcal{C}}$,
+Based on the discussion on the null eigenvectors in Appendix (C), we can find a null eigenvector $\mathcal{V}_{a}^{(\mathcal{E})}$ to build a new first-class constraint $\tilde{\mathcal{C}}$,
 $\int\mathrm{d}^{3}x\,\tilde{\mathcal{C}}(\vec{x})V(\vec{x})=\int\mathrm{d}^{3}x\left(\begin{array}[]{cccc}p&\pi&p^{kl}&\mathcal{C}^{\prime}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}{}_{1}\right)_{ij}^{(\mathcal{E})}&\mathcal{\mathcal{V}}{}_{2}^{(\mathcal{E})}&\mathcal{\mathcal{V}}{}_{3}^{(\mathcal{E})}\end{array}\right)(\vec{x})\approx 0.$ (117)
@@
 The number of DOFs is thus
-#$\#_{\text{dof}}$ $=\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (118)
+$\#_{\text{dof }}$ $=\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (118)
 […]
-=2=2.= 2 .
-Finally, let us discuss another 2DOF condition $\mathcal{S}_{2}^{(1)}$. By applying the consistency condition of the constraint $\mathcal{C}_{2}$, we can obtain the constraint $\Phi$,
+$=2.$
+Finally let us discuss another 2DOF condition $\mathcal{S}_{2}^{(1)}$. By applying the consistency condition of constraint $\mathcal{C}_{2}$, we can obtain the constraint $\Phi$,
 $\Phi(\vec{x})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),H_{C}\right]$ (119)
@@
 $\frac{\partial c_{3}}{\partial\phi}\mathrm{D}^{i}\mathrm{D}_{j}B_{kl}.$ (120)
-Therefore, requiring the 2DOF condition $\mathcal{S}_{2}^{(1)}=\det\mathcal{D}^{ab}\approx 0$ also implies that $\frac{\partial c_{3}}{\partial\phi}$ may vanish as one of the conditions for the matrix $\mathcal{D}^{ab}$ to be degenerate, which is sufficient to satisfy the 2DOF condition $\mathcal{S}_{2}^{(2)}$.
+So when we require the 2DOF condition $\mathcal{S}_{2}^{(1)}=\det\mathcal{D}^{ab}\approx 0$, it will also imply that $\frac{\partial c_{3}}{\partial\phi}$ may be vanishing as one of the conditions for the matrix $\mathcal{D}^{ab}$ to be degenerate, which is sufficient to satisfy 2DOF condition $\mathcal{S}_{2}^{(2)}$.
 VI Conclusion
 In this work, we investigated a class of spatially covariant gravity theories with a non-dynamical scalar field.
-We briefly described our model in Sec.
+We briefly describe our model in Sec.
 II.
-The general Lagrangian is given by (3), which has been shown to have two tensorial and one scalar DOFs [85].
-The purpose of this work is to determine the conditions on the Lagrangian, which we dub the 2DOF conditions, under which only two degrees of freedom propagate.
-A perturbative approach was taken in [86] to derive conditions such that the scalar mode is eliminated at linear order in perturbations.
-In this work, we employ a strict Hamiltonian constraint analysis to derive conditions such that only two DOFs propagate in the nonperturbative sense.
+The general Lagrangian is given by (3), which has been proved to have two tensorial and one scalar DOFs [85].
+The purpose of this work is to determine the conditions on the Lagrangian, which we dub the 2DOF conditions, under which only two degrees of freedom are propagating.
+A perturbative approach has been taken in [86] to derive conditions such that the scalar mode is eliminated at linear order in perturbations.
+In this work we employ the strict Hamiltonian constraint analysis to derive the conditions such that that only two DOFs are propagating in the nonperturbative sense.
 In Sec.
@@
 IV, by requiring the degeneracy of the Dirac matrix, we find that two conditions are required to make the theory propagate only 2DOFs.
-The first condition (38) states that the matrix $\mathcal{B}^{ab}$ given in (39) must be degenerate, which results in a secondary constraint.
-To fully eliminate the scalar mode, a second condition is necessary.
+The first condition (38) states that the matrix $\mathcal{B}^{ab}$ given in (39) must be degenerate, which will result in a secondary constraint.
+In order to fully eliminate the scalar mode, a second condition is necessary.
 The second condition can be divided into two categories according to their different effects on the Dirac matrix.
-In the first case, the condition (49) generally turns a second-class constraint into a first-class constraint.
-In the second case, the condition (50) generates another secondary constraint $\Phi$.
-We would like to emphasize that even if the Lagrangian satisfies both 2DOF conditions, the number of DOFs does not necessarily decrease.
-A special case is when the constraints are linearly dependent, which implies that the number of DOFs may remain unchanged.
-Another special case, which we refer to as the non-physical branch of the 2DOF conditions, renders the theory trivial or physically unacceptable (e.g., lacking tensor modes).
-One must therefore be careful when dealing with the 2DOF conditions in order to select the physical case, i.e., theories with precisely two degrees of freedom.
+In the first case, the condition (49) will generally turn a second-class constraint into a first-class constraint.
+In the second case, the condition (50) will generate another secondary constraint $\Phi$.
+We would like to emphasize that even the Lagrangian satisfies the both two 2DOF conditions, the number of DOFs do not necessarily decrease.
+A special case is that the constraints are linearly dependent, which implies that the number of DOFs may remain unchanged.
+Another special case, which we refer to as the non-physical branch of the 2DOF conditions, will make the theory trivial or physically unacceptable (e.g., without tensor modes).
+One thus must be careful when dealing with the 2DOF conditions in order to pick up the physical case, i.e., theories with precisely two degrees of freedom.
 In Sec.
-V, based on the spatially covariant monomials classified in [57], [58], we consider a concrete Lagrangian (60) built from monomials of $d=2$ (i.e., with two derivatives) as an illustration of our formal analysis.
-Since the Dirac matrix of this model contains derivatives of the delta function, it is complicated to calculate the matrix form of the 2DOF conditions for this concrete model.
-Instead, we choose an equivalent but more efficient method to obtain the 2DOF conditions, which is to find the null eigenvector $\mathcal{V}_{a}$ satisfying (40).
+V, based on the spatially covariant monomials classified in [57, 58], we consider a concrete Lagrangian (60) built of monomials of $d=2$ (i.e., with two derivatives) as an illustration of our formal analysis.
+Since the Dirac matrix of this model contains derivatives of the delta function, it is complicated to calculate the matrix form of the 2DOF conditions of this concrete model.
+Instead, we choose an equivalent but more efficient way to obtain the 2DOF conditions, which is to find out the null eigenvector $\mathcal{V}_{a}$ satisfing (40).
 The resulting 2DOF conditions are given in (66) and (111).
 In particular, for this concrete model, we are able to eliminate the non-physical branch and the linear constraints branch of the 2DOF conditions.
-Finally, we find a set of coefficients (114), and the resulting Lagrangian corresponds to a spatially covariant gravity theory with an auxiliary scalar field that propagates only two DOFs.
+Finally, we find a set of coefficients (114), and the resulting Lagrangian corresponds to a spatially covariant gravity theory with an auxiliary scalar field which propagates only two DOFs.
 Acknowledgements.
@@
 Appendix A Comparison with scalar-tensor theory in the spatial gauge
-The spatially covariant gravity with an auxiliary scalar field was firstly proposed in [85], where the Hamiltonian analysis was performed to show that the theory propagates 3 degrees of freedom.
+The spatially covariant gravity the an auxiliary scalar field was firstly proposed in [85], in which the Hamiltonian analysis was performed in order to show that the theory propagates 3 degrees of freedom.
 This idea was originally motivated by generally covariant scalar-tensor theory when the scalar field possesses a spacelike gradient.
-After choosing the so-called “spatial gauge” as in [85], the resulting action takes a form similar to (1).
-However, we should note that “spatially covariant gravity with an auxiliary scalar field” and “scalar-tensor theory in the spatial gauge” are completely different theories.
-First, we will show the difference between (1) and the scalar-tensor theory in the spatial gauge. Let us consider a general action of scalar-tensor theory
+After choosing the so-called “spatial gauge” as in [85], the resulting action takes the form that is similar to (1).
+However, we should notice that “spatially covariant gravity the an auxiliary scalar field” and “scalar-tensor theory in the spatial gauge” are completely different theories.
+First we will show the difference between (1) and the scalar-tensor theory in the spatial gauge. Let us consider a general action of scalar-tensor theory
 $S_{\mathrm{GST}}=\int\mathrm{d}^{4}x\sqrt{-g}\mathcal{L}\left(\phi;g_{ab},\varepsilon_{abcd},{}^{4}\!R_{abcd};\nabla_{a}\right),$ (121)
-where the Lagrangian involves a scalar field $\phi$, the spacetime metric $g_{ab}$, the spacetime curvature tensor ${}^{4}\!R_{abcd}$ as well as their covariant derivatives.
-The 4-dimensional Levi-Civita tensor $\varepsilon_{abcd}$ encodes possible parity violation effects.
-In a cosmological context, the scalar field $\phi$ is assumed to possess a timelike gradient so that the so-called unitary gauge with $\phi=t$ can be chosen.
-Here we consider the opposite situation by assuming that the scalar field possesses a spacelike gradient [^3] Similar ideas appear in “elastic inflation” [87] and “solid inflation” [88], where it is the Nambu-Goldstone boson breaking spatial diffeomorphism that plays the role of the scalar field..
-Contrary to the unitary gauge which is defined by requiring $\mathrm{D}_{a}\phi=0$, we now choose a gauge in which
-£$\pounds_{\bm{n}}\phi=0,$ (122)
+where the Lagrangian involves a scalar field $\phi$, spacetime metric $g_{ab}$, the spacetime curvature tensor $^{4}\!R_{abcd}$ as well as their covariant derivatives.
+The 4-dimension Levi-Civita tensor $\varepsilon_{abcd}$ encodes possible parity violation effects.
+In cosmological context, the scalar field $\phi$ is assumed to possesses a timelike gradient so that the so-called unitary gauge with $\phi=t$ can be chosen.
+Here we consider the contrary situation by assuming that the scalar field possesses a spacelike gradient [^3] Similar ideas appear in “elastic inflation” [87] and “solid inflation” [88] where it is the Nambu-Goldstone boson breaking spatial diffeomorphism that plays the role of the scalar field..
+Contrary to the unitary gauge which is defined by requiring $\mathrm{D}_{a}\phi=0$, we now can choose a gauge in which
+$\pounds_{\bm{n}}\phi=0,$ (122)
 where $n^{a}$ is the normal vector to the hypersurfaces.
-This can be understood as choosing hypersurfaces such that the normal vector $n^{a}$ exactly lies on the constant $\phi$ hypersurfaces.
-As a result, the value of $\phi$ does not change when transported along the normal vector.
-This choice of spatial hypersurfaces is referred to as the “spatial gauge” in [85].
-The normal vector $n^{a}$ is normalized by $n^{a}n_{a}=-1$. The induced metric on the spatial hypersurfaces is, as usual,
+This can be understood that we choose hypersurfaces such that the normal vector $n^{a}$ exactly lie on the the constant $\phi$ hypersurfaces.
+As a result, the value of $\phi$ does not change when being transported along the normal vector.
+This choice of spatial hypersurfaces is referred as “spatial gauge” in [85].
+The normal vector $n^{a}$ is normalized by $n^{a}n_{a}=-1$. The induced metric on the spatial hypersurfaces is as usual
 $h_{ab}=g_{ab}+n_{a}n_{b}.$ (123)
@@
 […]
-where $\mathrm{D}_{a}$ is the covariant derivative compatible with the induced metric $h_{ab}$.
-In the last equality of (124) we have used the fact that £$\pounds_{\bm{n}}\phi=0$ in the spatial gauge.
-This is exactly the opposite of the unitary gauge, where $\mathrm{D}_{a}\phi=0$ and thus $\nabla_{a}\phi\rightarrow-n_{a}\pounds_{\bm{n}}\phi$.
-After performing the 3+1 decomposition and choosing the spatial gauge, the action (121) can be written in the form
+where $\mathrm{D}_{a}$ is the covariant derivative compatible with induced metric $h_{ab}$.
+In the last equality of (124) we have used the fact that $\pounds_{\bm{n}}\phi=0$ in the spatial gauge.
+This is exactly contrary to what in the unitary gauge where $\mathrm{D}_{a}\phi=0$ and thus $\nabla_{a}\phi\rightarrow-n_{a}\pounds_{\bm{n}}\phi$.
+After making the 3+1 decomposition and choosing the spatial gauge, the action (121) can written in the form
 $S_{\mathrm{GST}}^{\mathrm{(s.g.)}}=\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij};\mathrm{D}_{i},\pounds_{\bm{n}}\right),$ (125)
 where “s.g.” stands for the spatial gauge and we have fixed the spatial coordinates adapted to the spacelike hypersurfaces.
-At first glance, the action (125) takes the same form as (1); however, the crucial difference is that in the spatial gauge, the scalar field $\phi$ can be viewed as a time-independent but space-dependent field, which breaks spatial diffeomorphism.
-In other words, $\phi=\phi(\bm{x})$ in (125) cannot be viewed as a dynamical nor auxiliary variable, but is actually a “function” of spatial coordinates with fixed values.
-This is completely different from (1), which has spatial covariance and where $\phi$ plays the role of an auxiliary field.
-Next, we show that the generally covariant version (i.e., the correspondence) of (1) is nothing but a “two-field” scalar-tensor theory.
+At the first glance, the action (125) takes the same form as (1), however, the crucial difference is that in the spatial gauge, the scalar field $\phi$ can be viewed as a time-independent but space-dependent field, which breaks spatial diffeomorphism.
+In other words, $\phi=\phi(\bm{x})$ in (125) cannot be viewed as a dynamical nor auxiliary variable, which is actually a “function” of space coordinates with fixed values.
+This is completely different from (1) which has spatial covariance, in which $\phi$ plays the role of an auxiliary field.
+Then we show that the generally covariant version (i.e., correspondence) of (1) is nothing but a “two-field” scalar-tensor theory.
 Let us take $\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi$ as an example.
@@
 $\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi=h^{ij}\mathrm{D}_{i}\phi\mathrm{D}_{j}\phi\rightarrow\left(g^{ab}+u^{a}u^{b}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (126)
-where $u_{a}=-N\nabla_{a}\Phi$ with $\Phi$ being the scalar field defining the spacelike hypersurfaces.
-In other words, the spacelike hypersurfaces are defined by $\Phi=\mathrm{const.}$.
+where $u_{a}=-N\nabla_{a}\Phi$ with $\Phi$ is the scalar field defining the spacelike hypersurfaces.
+In other words, the spacelike hypersurfaces are defined by hypersurfaces with $\Phi=\mathrm{const.}$.
 In (126), $N=1/\sqrt{-\left(\nabla\Phi\right)^{2}}$, which reduces to the lapse function when fixing the so-called unitary gauge with $\Phi=t$.
 By expanding (126) explicitly, we get
-$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi\rightarrow\left(g^{ab}-\frac{\nabla^{a}\Phi\nabla^{b}\Phi}{\left(\nabla\Phi\right)^{2}}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (127)
+$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi\rightarrow\left(g^{ab}-\frac{\nabla^{a}\Phi N\nabla^{b}\Phi}{\left(\nabla\Phi\right)^{2}}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (127)
 which is clearly a two-field scalar-tensor theory term.
 Appendix B Constraint algebra
-In this appendix, we present the explicit expressions for the Poisson brackets.
+In this appendix, we show the explicit expression for the Poisson brackets.
 The Poisson brackets among constraints are
@@
 $\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $\frac{\delta S}{\delta h_{ij}(\vec{x})}-\frac{1}{N(\vec{x})}\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{z})}N(\vec{z})B_{kl}(\vec{z}),$ (142)
-$\left[\mathcal{C}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $2B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}-\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta N(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}),$ (143)
+$\left[\mathcal{C}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $2B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}-\int\mathrm{d}^{3}z\frac{\delta S}{\delta N(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}),$ (143)
 $\left[\frac{\delta S}{\delta\phi(\vec{x})},H_{\mathrm{C}}\right]$ $=$ $\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}).$ (144)
@@
 If we wish to find the null eigenvector of the degenerate matrix $\mathcal{B}^{ab}$, one method is to calculate the non-zero rows (or columns) of its corresponding adjoint matrix $\mathcal{B}_{ab}^{*}$, and to multiply the corresponding elements by $(-1)^{a+b}$ to build the null eigenvector.
-For simplicity, we assume that $\mathcal{B}^{ab}$ is a simple $\times 3$ 3 matrix and the rows and columns correspond to constraints $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$.
+For simplicity, we assume that $\mathcal{B}^{ab}$ is a simple $3\times 3$ matrix and the rows and columns correspond to constraints $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$.
 To verify that this form of null eigenvector is feasible, we assume that the matrix $\mathcal{B}^{ab}$ is degenerate and without loss of generality we take the first row of its adjoint matrix $\mathcal{B}_{ab}^{*}$ to form a null eigenvector
@@
 $=$ $\int\mathrm{d}^{3}z\left(\begin{array}[]{ccc}\mathcal{B}^{1i}&\mathcal{B}^{2i}&\mathcal{B}^{3i}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)^{T}(\vec{z},\vec{y})$ (171)
-$=$ 0 0.0 . (172)
+$=$ $0.$ (172)
 In practical calculations, in order to avoid integrating the delta function of higher order, we need to multiply the null eigenvector by a greatest common divisor, i.e.,